### v1

# Nuclear Toeplitz operators between Fock spaces

Abstract.
We study Toeplitz operators with measure-valued symbols acting between Fock spaces. Given $1\leq p,q\leq\infty$ and a Borel measure $\mu$ on $\mathbb{C}$, we investigate when the associated Toeplitz operator
$T_{\mu}:F^{p}_{\alpha}\to F^{q}_{\alpha}$
belongs to the nuclear class. For positive measures $\mu$ and in the range $1\leq q\leq p\leq\infty$, we obtain necessary and sufficient conditions for the nuclearity of $T_{\mu}$ in terms of the Berezin transform of $\mu$. As a consequence, nuclearity in this setting exhibits a rigidity property: if $T_{\mu}$ is nuclear from $F^{p}_{\alpha}$ to $F^{q}_{\alpha}$ for some $q\leq p$, then it is nuclear for all such $q$. In the case $p<q$, we show that the situation is more delicate. We provide separate necessary and sufficient conditions for nuclearity, indicating that the Berezin transform alone does not yield a complete characterization. The proofs rely on tools from Banach space operator theory combined with kernel estimates on Fock spaces. Our results extend naturally to Fock spaces on $\mathbb{C}^{n}$.
Key words and phrases:
Nuclear operator, Toeplitz operator, Fock space
1991 Mathematics Subject Classification:
Primary 47B10 Secondary 47B35
*Corresponding author.
1. Introduction
Let $\alpha>0$ and $p>0$, the notation $L_{\alpha}^{p}$ refers to the space of all Lebesgue measurable functions $f$ on $\mathbb{C}$ for which $f(z)\mathrm{e}^{-\alpha|z|^{2}/2}\in L^{p}(\mathbb{C},\mathrm{d}A)$, where $\mathrm{d}A$ is the Lebesgue area measure on $\mathbb{C}$. For $f\in L_{\alpha}^{p}$, its norm is given by
$\|f\|_{\alpha,p}^{p}=\frac{p\alpha}{2\pi}\int_{\mathbb{C}}\left|f(z)\mathrm{e}^{-\frac{\alpha}{2}|z|^{2}}\right|^{p}\,dA(z).$
Correspondingly, for $\alpha>0$ and $p=\infty$, we denote by $L_{\alpha}^{\infty}$ the space of measurable functions $f$ such that
$\|f\|_{\infty,\alpha}=\operatorname{ess\,sup}\left\{|f(z)|\mathrm{e}^{-\alpha|z|^{2}/2}:z\in\mathbb{C}\right\}<\infty.$
The Fock space $F_{\alpha}^{p}$ is defined as the subspace of $L_{\alpha}^{p}$ consisting of entire functions. It is straightforward to verify that $F_{\alpha}^{p}$ becomes a Banach space when $1\leq p\leq\infty$. Let $f_{\alpha}^{\infty}$ be the closed subspace of $F_{\alpha}^{\infty}$ consisting of $f$ satisfying
$\lim_{|z|\to\infty}f(z)e^{-\frac{\alpha}{2}|z|^{2}}=0.$
When $p=2$, the space $F_{\alpha}^{2}$ is a Hilbert space equipped with the inner product
$\langle f,g\rangle_{F_{\alpha}^{2}}=\frac{\alpha}{\pi}\int_{\mathbb{C}}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}\,dA(z).$
As a reproducing kernel Hilbert space, its reproducing kernel is $K_{\alpha}(z,w)=e^{\alpha z\overline{w}}$. The normalized reproducing kernel at $z$ is denoted by $k_{z}(w)=e^{\alpha\overline{z}w-\frac{\alpha}{2}|z|^{2}}$, which is a unit vector in $F_{\alpha}^{p}$ for every $0<p\leq\infty$.
For $1\leq p<\infty$, let $p^{\prime}$ be the conjugate exponent satisfying $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. We can identify the dual of $F_{\alpha}^{p}$ with $F_{\alpha}^{p^{\prime}}$, and the dual of $f_{\alpha}^{\infty}$ with $F_{\alpha}^{1}$, via the integral pairing
$\langle f,g\rangle_{F_{\alpha}^{2}}=\frac{\alpha}{\pi}\int_{\mathbb{C}}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}\,dA(z),$
and under this identification, the norms on the dual spaces and the Fock spaces are equivalent (see [19, Corollary 2.25, Theorem 2.26]).
Given $\alpha>0$, consider the Gaussian measure
$d\lambda_{\alpha}(z)=\frac{\alpha}{\pi}e^{-\alpha|z|^{2}}\,dA(z).$
Define the integral operator $P_{\alpha}$ on $L_{\alpha}^{p}$ by
$P_{\alpha}f(z)=\int_{\mathbb{C}}K_{\alpha}(z,w)f(w)\,d\lambda_{\alpha}(w),\quad f\in L_{\alpha}^{p}.$
It is known that $P_{\alpha}:L_{\alpha}^{p}\to F_{\alpha}^{p}$ is bounded for all $p$, and it acts as the identity on $F_{\alpha}^{p}$, that is, $P_{\alpha}f=f$ for $f\in F_{\alpha}^{p}$ (see [19, Corollary 2.22]).
For a function $\varphi\in L^{\infty}(\mathbb{C})$, the associated Toeplitz operator $T_{\varphi}:F_{\alpha}^{2}\to F_{\alpha}^{2}$ is defined by
$T_{\varphi}(f)=P_{\alpha}(\varphi f),\quad f\in F_{\alpha}^{2}.$
More generally, if $\mu$ is a complex Borel measure on $\mathbb{C}$, the Toeplitz operator $T_{\mu}$ is given by
$T_{\mu}(f)(z)=\frac{\alpha}{\pi}\int_{\mathbb{C}}K_{\alpha}(z,w)f(w)\mathrm{e}^{-\alpha|w|^{2}}\,d\mu(w),\quad z\in\mathbb{C},$
provided that the integrability condition
$\int_{\mathbb{C}}|K_{\alpha}(z,w)|\mathrm{e}^{-\alpha|w|^{2}}\,d|\mu|(w)<\infty$ (1.1)
holds for every $z\in\mathbb{C}$. Due to the exponential nature of the kernel, condition (1.1) is equivalent to
$\int_{\mathbb{C}}|K_{\alpha}(z,w)|^{2}\mathrm{e}^{-\alpha|w|^{2}}\,d|\mu|(w)<\infty,\quad z\in\mathbb{C}.$ (1.2)
Let
$K=\operatorname{span}\{K_{z}:z\in\mathbb{C}\}.$ (1.3)
The set $K$ is dense in $F_{\alpha}^{p}$ for every $1\leq p<\infty$ and also in $f_{\alpha}^{\infty}$. Under assumption (1.2), the operator $T_{\mu}$ is well-defined on $K$ by the Cauchy-Schwarz inequality, and hence on a dense subset of each $F_{\alpha}^{p}$ ($1\leq p<\infty$) and $f_{\alpha}^{\infty}$. For a measure $\mu$ satisfying (1.1), its Berezin transform is defined as
$\widetilde{\mu}(z)=\frac{\alpha}{\pi}\int_{\mathbb{C}}|k_{z}(w)|^{2}\mathrm{e}^{-\alpha|w|^{2}}\,d\mu(w)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\mathrm{e}^{-\alpha|z-w|^{2}}\,d\mu(w).$ (1.4)
Toeplitz operators on Fock spaces form an important class of operators connecting complex analysis, operator theory, and the theory of Banach function spaces. Owing to the reproducing kernel structure of Fock spaces, Toeplitz operators provide a natural framework for studying how analytic and measure-theoretic properties of symbols are reflected in operator-theoretic behavior. The theory of Toeplitz operators on Fock spaces has been extensively developed in the Hilbert space setting. For the classical Fock space $F^{2}_{\alpha}$, Toeplitz operators with bounded or measure-valued symbols have been studied from various perspectives, including boundedness, compactness, and Schatten class membership, see references [1, 2, 3, 9]. In particular, Isralowitz and Zhu [13] obtained complete characterizations of boundedness, compactness, and Schatten class properties of Toeplitz operators with positive Borel measures as symbols.
Beyond the Hilbertian framework, Hu and Lv [11, 12] systematically investigated Toeplitz operators acting between different Fock spaces $F^{p}_{\alpha}$ and $F^{q}_{\alpha}$ for $1<p,q<\infty$. They characterized boundedness and compactness of Toeplitz operators with positive Borel measure symbols in terms of averaging functions and Berezin-type transforms. A notable feature of their results is a rigidity phenomenon specific to Fock spaces: boundedness (or compactness) from $F^{p}_{\alpha}$ to $F^{q}_{\alpha}$ for some $p\leq q$ automatically implies boundedness (or compactness) for all such $q$. This behavior differs substantially from what is known in the Bergman space setting. The groundbreaking work of Hu and Lv has stimulated extensive scholarly investigations into Toeplitz and Hankel operators across diverse weighted Fock spaces [16, 19].
In contrast to the well-developed theory of bounded and compact Toeplitz operators, considerably less is known about Toeplitz operators belonging to finer operator ideals on Fock spaces, especially in the non-Hilbertian case. Among these ideals, the class of nuclear operators plays a distinguished role in Banach space operator theory. Nuclearity represents a strong form of compactness and depends sensitively on the geometry of the underlying Banach spaces. As a result, techniques based on Hilbert space methods or Schatten class theory are generally insufficient to address nuclearity questions when $p\neq 2$.
The aim of this paper is to study nuclear Toeplitz operators acting between Fock spaces. Given $1\leq p,q\leq\infty$ and a Borel measure $\mu$ on $\mathbb{C}$, we investigate conditions under which the Toeplitz operator
$T_{\mu}:F^{p}_{\alpha}\longrightarrow F^{q}_{\alpha}$
belongs to the nuclear class $N(F^{p}_{\alpha},F^{q}_{\alpha})$. Our focus is primarily on Toeplitz operators with positive symbols, although several results extend to complex measures under suitable assumptions.
Our main results provide a detailed description of nuclear Toeplitz operators in different parameter regimes. When $1\leq q\leq p\leq\infty$ and $\mu\geq 0$, we obtain necessary and sufficient conditions for $T_{\mu}$ to be nuclear in terms of the integrability of the Berezin transform of $\mu$ (Theorem 3.7). As a consequence, we show that nuclearity in this range exhibits a rigidity property analogous to that observed for boundedness: if $T_{\mu}$ is nuclear from $F^{p}_{\alpha}$ to $F^{q}_{\alpha}$ for some $q\leq p$, then it is nuclear for all such $q$.
When $p<q$, the situation becomes more subtle. In this case, we establish separate necessary and sufficient conditions for nuclearity, showing that the Berezin transform alone does not yield a complete characterization. This reflects an inherent asymmetry between the domain and target spaces in the nuclear setting and distinguishes the nuclear theory from previously studied mapping properties of Toeplitz operators on Fock spaces.
The proofs rely on tools from Banach space operator theory, including duality properties of nuclear operators, approximation and Radon–Nikodým properties of Fock spaces, and precise kernel estimates associated with the Berezin transform. The arguments presented here extend naturally to Fock spaces on $\mathbb{C}^{n}$, and the main results remain valid in higher dimensions without essential changes.
The paper is organized as follows. In Section 2, we recall basic definitions and preliminary results concerning nuclear operators and Fock spaces. Section 3 is devoted to the characterization of nuclear Toeplitz operators between Fock spaces and contains our main results. In Section 4, we prove density results showing that Toeplitz operators with compactly supported continuous symbols are dense in the space of nuclear operators between Fock spaces with respect to the nuclear norm. As a consequence, we recover and extend known results on trace ideals in the Hilbert space case $p=2$.
2. Preliminaries
This section collects known results and establishes basic facts concerning nuclear operators, which will be used in the subsequent sections.
Let $X$ and $Y$ be complex Banach spaces. The Banach space consisting of all bounded linear operators from $X$ to $Y$ is represented by $\mathcal{L}(X,Y)$. The dual space of a Banach space $X$ is denoted by $X^{\prime}$. For $x\in X$ and $x^{\prime}\in X^{\prime}$, we write $\langle x,x^{\prime}\rangle$ for the dual pairing $x^{\prime}(x)$.We say that an operator $T\in\mathcal{L}(X,Y)$ is nuclear provided that it admits a representation of the form
$T=\sum_{j=1}^{\infty}x_{j}^{\prime}\otimes y_{j},$ (2.1)
for sequences $(x_{j}^{\prime})\subset X^{\prime}$ and $(y_{j})\subset Y$ satisfying $\sum_{j=1}^{\infty}\|x_{j}^{\prime}\|_{X^{\prime}}\|y_{j}\|_{Y}<\infty$. Let $x^{\prime}\in X^{\prime},y\in Y$ , we use the notation $x^{\prime}\otimes y$ for the rank-one operator from $X$ to $Y$ defined by
$x^{\prime}\otimes y(x)=\langle x,x^{\prime}\rangle y,\quad x\in X.$
The nuclear norm of $T$ is given by
$\|T\|_{\mathcal{N}}:=\inf\sum_{j=1}^{\infty}\|x_{j}^{\prime}\|_{X^{\prime}}\|y_{j}\|_{Y},$
The infimum in this definition ranges over all representations of the type (2.1). we write $\mathcal{N}(X,Y)$ for the class of all nuclear operators mapping $X$ into $Y$. Endowed with the norm $\|\cdot\|_{\mathcal{N}}$, it becomes a Banach space, and the following estimate holds for any $A\in\mathcal{L}(X_{0},X)$, $T\in\mathcal{N}(X,Y)$, and $B\in\mathcal{L}(Y,Y_{0})$:
$\|BTA\|_{\mathcal{N}}\leq\|A\|\,\|T\|_{\mathcal{N}}\,\|B\|.$
In the case where $E$ and $F$ coincide, we write $\mathcal{N}(E)$ for $\mathcal{N}(E,E)$. We say that a Banach space $E$ has the approximation property provided that for any compact set $K\subset E$ and any $\epsilon>0$, one can find a finite-rank operator $L:E\to E$ satisfying
$\|x-Lx\|<\epsilon\quad\text{for all }x\in K.$
For a Banach space $E$ with the approximation property, the nuclear trace of an operator $T\in\mathcal{N}(E)$ can be defined by
$\operatorname{Tr}_{E}(T)=\sum_{j=1}^{\infty}\langle x_{j},x_{j}^{\prime}\rangle,$
where $T=\sum_{j=1}^{\infty}x_{j}^{\prime}\otimes x_{j}$ is any representation as in (2.1). It follows from [10] Theorem V.1.2 that the value of the trace does not depend on the particular representation chosen. Standard references for the general theory of operator ideals include [5, 6, 10, 17].
Definition 2.1.
Let $X$ be a Banach space, and let $(\Omega,\Sigma,\mu)$ be any finite measure space, if for every $\mu$-continuous vector measure $G:\Sigma\to X$ of bounded variation, we can find a function $g\in L^{1}(\mu,X)$ for which
$G(E)=\int_{E}g\,d\mu\qquad\text{holds for all }E\in\Sigma,$
we say $X$ has the Radon-Nikodým property.
For each nonnegative integer $n$, define
$e_{n}(z)=\sqrt{\frac{\alpha^{n}}{n!}}z^{n}.$
It is well known that $\{e_{n}\}_{n=0}^{\infty}$ forms an orthonormal basis for $F_{\alpha}^{2}$. For a general exponent $1<p<\infty$, the same system $\{e_{n}\}_{n=0}^{\infty}$ is also a Schauder basis for $F_{\alpha}^{p}$, as can be seen by expanding each $f\in F_{\alpha}^{p}$ in a Taylor series about the origin.
Proposition 2.2.
For $1<p<\infty$, the Fock space $F_{\alpha}^{p}$ is a Banach space that enjoys both the approximation property and the Radon–Nikodým property.
The approximation property follows from the fact that $F_{\alpha}^{p}$ has a Schauder basis (see [9, Proposition 2.1]). The Radon–Nikodým property holds because $F_{\alpha}^{p}$ is reflexive (see [7, Corollary 3.13]). ∎
Let $X$ be a Banach space, if $x\in X$, define $x^{\prime\prime}:X^{\prime}\rightarrow\mathbb{C}$ by
$x^{\prime\prime}(x^{\prime})=x^{\prime}(x)$
for every $x^{\prime}\in X^{\prime}$. The map $\tau:X\rightarrow X^{\prime\prime}$ defined by $\tau(x)=x^{\prime\prime}$ for all $x\in X$ is called the canonical map. Recall that $\tau$ is an isometry from $X$ onto $X^{\prime\prime}$ when $X$ is reflexive.
Theorem 2.3.
Let X and Y be Banach spaces such that Y is reflexive. The adjoint map $T\rightarrow T^{*}$ is an isometry from $\mathcal{N}(X,Y)$ onto $\mathcal{N}(Y^{\prime},X^{\prime})$.
Let $T=\sum_{j=1}^{\infty}x_{j}^{\prime}\otimes y_{j}\in\mathcal{N}(X,Y)$, where $(x_{j}^{\prime})\in X^{\prime},(y_{j})\in Y$. For any $y^{\prime}\in Y^{{}^{\prime}},x\in X$, we have
$\displaystyle T^{*}(y^{\prime})(x)$ $\displaystyle=y^{\prime}(Tx)$
$\displaystyle=y^{\prime}(\sum_{j=1}^{\infty}x_{j}^{\prime}(x)y_{j})$
$\displaystyle=\sum_{j=1}^{\infty}x_{j}^{\prime}(x)y^{\prime}(y_{j})$
$\displaystyle=\sum_{j=1}^{\infty}y_{j}^{\prime\prime}(y^{\prime})x_{j}^{\prime}(x).$
Then $T^{*}=\sum_{j=1}^{\infty}y_{j}^{\prime\prime}\otimes x_{j}^{\prime}$, and $\|T^{*}\|_{\mathcal{N}}\leq\sum_{j=1}^{\infty}\|y_{j}^{\prime\prime}\|_{Y^{\prime\prime}}\|x_{j}^{\prime}\|_{X^{\prime}}=\sum_{j=1}^{\infty}\|{y_{j}\|_{Y}\|x_{j}^{\prime}\|_{X^{\prime}}}$, the inequality $\|T^{*}\|_{\mathcal{N}}\leq\|T\|_{\mathcal{N}}$ follows by taking the infimum over all representations of the form (2.1). On the other hand, let $K\in\mathcal{N}(Y^{\prime},X^{\prime})$, since $Y$ is reflexive, there are sequences $(a_{j}^{\prime})\in X^{\prime}$, $(b_{j})\in Y$ such that $K=\sum_{j=1}^{\infty}b_{j}^{\prime\prime}\otimes a_{j}^{\prime}$, let $L=\sum_{j=1}^{\infty}a_{j}^{\prime}\otimes b_{j}\in\mathcal{N}(X,Y)$, It is easy to check that $L^{*}=K$ and $\|L\|_{\mathcal{N}}\leq\|K\|_{\mathcal{N}}$. This prove the desire result. ∎
For the dual space of compact operator and nuclear operator, we have the following characterization.
Proposition 2.4.
([5, Proposition 16.7]) Let $E$ and $F$ be Banach spaces such that either $E^{\prime\prime}$ or $F^{\prime}$ has the approximation property, and either $E^{\prime\prime}$ or $F^{\prime}$ has the Radon–Nikodým property. Then the following isometric identities hold:
$\displaystyle(\mathcal{K}(E,F))^{\prime}$ $\displaystyle=\mathcal{N}\left(E^{\prime},F^{\prime}\right),$
$\displaystyle(\mathcal{K}(E,F))^{\prime\prime}$ $\displaystyle=\left(\mathcal{N}\left(E^{\prime},F^{\prime}\right)\right)^{\prime}=\mathcal{L}\left(E^{\prime\prime},F^{\prime\prime}\right),$
where the duality pairings are given by
$\displaystyle\langle S,T\rangle_{\mathcal{K},\mathcal{N}}$ $\displaystyle=\sum_{n=1}^{\infty}\left\langle S^{*}y_{n}^{\prime},x_{n}^{\prime\prime}\right\rangle_{E^{\prime},E^{\prime\prime}},$
$\displaystyle\langle T,S\rangle_{\mathcal{N},\mathcal{L}}$ $\displaystyle=\sum_{n=1}^{\infty}\left\langle y_{n}^{\prime},Sx_{n}^{\prime\prime}\right\rangle_{E^{\prime},E^{\prime\prime}},$
for $T=\sum_{n=1}^{\infty}x_{n}^{\prime\prime}\otimes y_{n}^{\prime}\in\mathcal{N}\left(E^{\prime},F^{\prime}\right)$.
Proposition 2.5.
Let $1<p,q<\infty$, then
$\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})^{\prime}=\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p}),\quad\mathcal{K}(F_{\alpha}^{p},F_{\alpha}^{q})^{\prime}=\mathcal{N}(F_{\alpha}^{q},F_{\alpha}^{p})$
where the duality is induced by the trace map:
$\displaystyle\langle S,T\rangle$ $\displaystyle=Tr(ST)=Tr(TS)$
$\displaystyle\langle A,B\rangle$ $\displaystyle=Tr(AB)=Tr(BA)$
where $S\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$, $T\in\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p})$, $A\in\mathcal{K}(F_{\alpha}^{p},F_{\alpha}^{q})$, $B\in\mathcal{N}(F_{\alpha}^{q},F_{\alpha}^{p})$.
We use $E\cong F$ to denote that the two Banach spaces $E$ and $F$ are isometrically isomorphic. Then, we have
$\displaystyle\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})^{\prime}$ $\displaystyle\cong\mathcal{N}((F_{\alpha}^{q})^{\prime},(F_{\alpha}^{p})^{\prime})^{\prime}$
$\displaystyle\cong\mathcal{L}((F_{\alpha}^{q})^{\prime\prime},(F_{\alpha}^{p})^{\prime\prime})$
$\displaystyle\cong\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p}).$
The first step is obtained from Theorem 2.3, the second step from Proposition 2.4, the last step follows from the reflexivity.
Let $S=\sum_{j=1}^{\infty}y_{j}\otimes x_{j}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ and $T\in\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p})$, applying Proposition 2.4, we obtain the duality bracket is given by
$\displaystyle\langle S,T\rangle$ $\displaystyle=\langle S^{*},\tau T\tau^{-1}\rangle_{\mathcal{N},\mathcal{L}}$
$\displaystyle=\sum_{j=1}^{\infty}\langle y_{j},\tau Tx_{j}\rangle$
$\displaystyle=\sum_{j=1}^{\infty}\langle Tx_{j},y_{j}\rangle$
$\displaystyle=Tr(TS)$
$\displaystyle=Tr(ST).$
The other proof is similar. ∎
For convenience, in the following sections, we write $A\lesssim B$ to mean that $A\leq CB$ for some positive constant $C$, whose exact value is insignificant. The relation $A\simeq B$ signifies that $A\lesssim B\lesssim A$.
3. Nuclear class Toeplitz operator
In this section, give a complex Borel measure $\mu$, we are going to characterize when the Toeplitz operator $T_{\mu}$ belongs to nuclear class.
Let $1\leq p\leq\infty$ and $p^{\prime}$ to be the conjugate exponent of $p$. For $f\in F_{\alpha}^{p^{\prime}}$, we regard $f$ as an element of $(F_{\alpha}^{p})^{\prime}$ under the pairing $\langle\cdot,\cdot\rangle_{F_{\alpha}^{2}}$, that is
$f(g)=\langle g,f\rangle_{F_{\alpha}^{2}}$
for all $g\in F_{\alpha}^{p}$. Then we have $\|f\|_{(F_{\alpha}^{p})^{\prime}}\simeq\|f\|_{\alpha,p^{\prime}}$.
Definition 3.1.
Let $T:F_{\alpha}^{p}\xrightarrow{}F_{\alpha}^{q}$ be a linear operator, where $1\leq p,q\leq\infty$, we define the Berezin transform of $T$ as
$\widetilde{T}(z)=\left\langle Tk_{z},k_{z}\right\rangle_{F_{\alpha}^{2}},\quad z\in\mathbb{C}.$
Theorem 3.2.
Let $1\leq p<\infty$, $1\leq q,s\leq\infty$. We have:
  1. (1)
    If $T\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{s})$, then $\widetilde{T}(z)\in L^{r}(dA)$ for all $1\leq r\leq\infty$.
  2. (2)
    If $T\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$, and $p<q$, then $\widetilde{T}(z)\in L^{r}(dA)$ for all $(\frac{1}{p^{\prime}}+\frac{1}{q})^{-1}\leq r\leq\infty$.
  3. (3)
    If $T\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$, and $q\leq p$, then $\widetilde{T}(z)\in L^{r}(dA)$ for all $1\leq r\leq\infty$.
Moreover,
$\|\widetilde{T}\|_{L^{r}}\lesssim\|T\|_{\mathcal{N}}.$
To prove (1), let $T=\sum_{j=1}^{\infty}f_{j}\otimes g_{j}$, where $\left(g_{j}\right)\subset F_{\alpha}^{s},\left(f_{j}\right)\subset F_{\alpha}^{1}$ be a nuclear representation. Hence
$\displaystyle\widetilde{T}(z)$ $\displaystyle=\langle Tk_{z},k_{z}\rangle_{F_{\alpha}^{2}}$
$\displaystyle=\sum_{j=1}^{\infty}\langle k_{z},f_{j}\rangle_{F_{\alpha}^{2}}\langle g_{j},k_{z}\rangle_{F_{\alpha}^{2}}$
$\displaystyle=\sum_{j=1}^{\infty}\overline{f_{j}(z)}g_{j}(z)e^{-\alpha|z|^{2}}.$
Let $1\leq r\leq\infty$, we can find $m\geq 1,n\geq s$ such that $\frac{1}{m}+\frac{1}{n}=\frac{1}{r}$. It follows from [19, Theorem 2.10], we have $f_{j}\in F_{\alpha}^{1}\subset F_{\alpha}^{m}$, $g_{j}\in F_{\alpha}^{s}\subset F_{\alpha}^{n}$, and
$\|f_{j}\|_{\alpha,m}\lesssim\|f_{j}\|_{\alpha,1},\|g_{j}\|_{\alpha,n}\lesssim\|g_{j}\|_{\alpha,s}.$
Then using Hölder’s inequality, we have
$\displaystyle\|\widetilde{T}\|_{L^{r}}$ $\displaystyle\leq\sum_{j=1}^{\infty}\|{f_{j}(z)}e^{-\frac{\alpha}{2}|z|^{2}}\|_{L^{m}}\|g_{j}(z)e^{-\frac{\alpha}{2}|z|^{2}}\|_{L^{n}}$
$\displaystyle\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,m}\|g_{j}\|_{\alpha,n}$
$\displaystyle\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,1}\|g_{j}\|_{\alpha,s}$
$\displaystyle\lesssim\sum_{j=1}^{\infty}\|{f_{j}}\|_{(f_{\alpha}^{\infty})^{\prime}}\|g_{j}\|_{\alpha,s}.$
Now, considering the infimum over all possible representations (2.1) yields
$\|\widetilde{T}\|_{L^{r}}\lesssim\|T\|_{\mathcal{N}}<\infty.$
The proof of (2) and (3) is similar. ∎
Lemma 3.3.
Let $1\leq p\leq\infty$, $z\rightarrow k_{z}$ is norm-continuous on $F_{\alpha}^{p}$.
Fix $z_{0}\in\mathbb{C}$, suppose $|z-z_{0}|<1$ and $1\leq p<\infty$. Then
$\|k_{z}-k_{z_{0}}\|_{\alpha,p}^{p}=\frac{\alpha p}{2\pi}\int_{\mathbb{C}}|(k_{z}(w)-k_{z_{0}}(w))e^{-\frac{\alpha}{2}|w|^{2}}|^{p}dA(w).$
Since $|z-z_{0}|<1$, there exists a constant $C>0$ such that $|(k_{z}(w)-k_{z_{0}}(w))e^{-\frac{\alpha}{2}|w|^{2}}|\leq Ce^{-\frac{\alpha}{4}|w|^{2}}$. According to the dominated convergence theorem, the conclusion holds for $1\leq p<\infty$. When $p=\infty$, the conclusion follows from the fact $\|f\|_{\alpha,\infty}\leq\|f\|_{\alpha,p}$ for any $1\leq p<\infty$. ∎
Lemma 3.4.
Let X be a Banach space, and let $\sum_{n=1}^{\infty}x_{n}$ be a absolutely convergent series in X with sum A, Then any rearrangement $\sum_{n=1}^{\infty}x_{\pi(n)}$ induced by a permutation $\pi:\mathbb{N}\rightarrow{\mathbb{N}}$ also converges to A.
Fix $\epsilon>0$. Choose $N\in\mathbb{N}$ such that:
$\sum_{k=N+1}^{\infty}\|x_{k}\|<\frac{\epsilon}{2}.$
Let $S_{n}=\sum_{k=1}^{n}x_{k}$. There exists $N_{1}\geq N$ such that for all $m\geq N_{1}$:
$\left\|\sum_{n=1}^{m}x_{n}-A\right\|<\frac{\epsilon}{2}.$
Choose $M\in\mathbb{N}$ such that $\{1,\ldots,N_{1}\}\subseteq\{\pi(1),\ldots,\pi(M)\}$. For any $m>M$, Split the rearranged partial sum:
$S^{\prime}_{m}=\sum_{\begin{subarray}{c}k\leq m\\ \pi(k)\leq N_{1}\end{subarray}}x_{\pi(k)}+\sum_{\begin{subarray}{c}k\leq m\\ \pi(k)>N_{1}\end{subarray}}x_{\pi(k)}.$
By the triangle inequality:
$\|S^{\prime}_{m}-A\|\leq\|S_{N_{1}}-A\|+\sum_{k=N_{1}+1}^{\infty}\|x_{k}\|<\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon.$
since $\epsilon$ is arbitrary. The rearranged series converges to $A$. ∎
For a fixed positive number $r$, let
$r\mathbb{Z}^{2}=\{nr+\mathrm{i}mr:n,m\in\mathbb{Z}\}$
Then, we define
$S_{r}=\{z=x+\mathrm{i}y:-r/2\leq x<r/2,\;-r/2\leq y<r/2\}.$
Clearly, we have the disjoint decomposition for $\mathbb{C}$:
$\mathbb{C}=\bigcup_{w\in r\mathbb{Z}^{2}}(S_{r}+w),$
Consequently, for every $f\in L^{1}(\mathbb{C},\mathrm{d}\mu)$ we have
$\int_{\mathbb{C}}f(z)\,\mathrm{d}\mu(z)=\sum_{w\in r\mathbb{Z}^{2}}\;\int_{S_{r}+w}f(z)\,\mathrm{d}\mu(z).$
Finally, let $\{a_{n}\}$ denote any fixed enumeration of the square lattice $r\mathbb{Z}^{2}$ as a sequence.
Theorem 3.5.
Let $\mu$ be a complex Borel measure on $\mathbb{C}$, and $|\mu|(\mathbb{C})<\infty$. Then $T_{\mu}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$. Moreover, we have
$\|T_{\mu}\|_{\mathcal{N}}\leq\frac{\alpha}{\pi}|\mu|(\mathbb{C}).$
Let $z\in\mathbb{C}$, we consider the operator $k_{z}\otimes k_{z}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$, i.e.
$(k_{z}\otimes k_{z})f=\langle f,k_{z}\rangle_{F_{\alpha}^{2}}k_{z}.$
It is obvious $\|k_{z}\otimes k_{z}\|_{\mathcal{N}}\leq 1$ for all $z\in\mathbb{C}$. For every $r>0$, note that
$\sum_{k=1}^{\infty}|\mu(S_{r}+a_{k})|\leq\sum_{k=1}^{\infty}|\mu|(S_{r}+a_{k})=|\mu|(\mathbb{C})<\infty,$
thus we can define
$T_{r}=\sum_{k=1}^{\infty}\mu(S_{r}+a_{k})k_{a_{k}}\otimes k_{a_{k}}\in\mathcal{N}{(f_{\alpha}^{\infty},F_{\alpha}^{1})}.$
Let $r_{n}=\frac{1}{2^{n}}$, fix any $\epsilon>0$, since $|\mu|(\mathbb{C})<\infty$, there exists $R>0$ such that
$|\mu|(\mathbb{C}\setminus\overline{B(0,R)}<\epsilon.$
According to Lemma 3.3, there exists a $\delta>0$ such that
$\|k_{z}-k_{w}\|_{\alpha,1}<\epsilon$
whenever $|z-w|<\delta$, and $z,w\in\overline{B(0,R+1)}$.
[…]
On the one hand, there exists an interger $N_{1}$ such that $r_{N}<\sqrt{2}\delta$ for all $N>N_{1}$. Fix a $N>max(N_{0},N_{1})$, then by Lemma 3.4, for any $N^{\prime}>N$, we have
$\displaystyle T_{r_{N}}-T_{r_{N^{\prime}}}=\sum_{k=1}^{\infty}\mu(S_{r_{N}}+a_{k}^{(N)})k_{a_{k}^{(N)}}\otimes k_{a_{k}^{(N)}}-\sum_{j=1}^{\infty}\mu(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})k_{a_{j}^{(N^{\prime})}}\otimes k_{a_{j}^{(N^{\prime})}}$
$\displaystyle=\sum_{k=1}^{\infty}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})](k_{a_{k}^{(N)}}\otimes k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}}\otimes k_{a_{j}^{(N^{\prime})}})$
$\displaystyle=I_{1}+I_{2}$
where
$\displaystyle I_{1}=\sum_{k=1}^{\infty}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})](k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\otimes k_{a_{j}^{(N^{\prime})}}$
$\displaystyle I_{2}=\sum_{k=1}^{\infty}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]k_{a_{k}^{(N)}}\otimes(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}}).$
Let
$S_{1}=\{k\in\mathbb{N},(S_{r_{N}}+a_{k}^{(N)})\cap\overline{B(0,R)}\neq\emptyset\},$
$S_{2}=\{k\in\mathbb{N},(S_{r_{N}}+a_{k}^{(N)})\cap\overline{B(0,R)}=\emptyset\}.$
It is easy to see that if $(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})\neq\emptyset,$ then we have $a_{k}^{(N)},$ $a_{j}^{(N^{\prime})}\in\overline{(S_{r_{N}}+a_{k}^{(N)})}$, and $|a_{k}^{(N)}-a_{j}^{(N^{\prime})}|<\delta$. Thus
$\displaystyle\|\sum_{k\in S_{1}}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]k_{a_{k}^{(N)}}\otimes(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\|_{\mathcal{N}}$
$\displaystyle\leq\sum_{k\in S_{1}}\sum_{j=1}^{\infty}|\mu|[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]\|(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\|_{\alpha,1}\|k_{a_{j}^{(N^{\prime})}}\|_{(f_{\alpha}^{\infty})^{\prime}}$
$\displaystyle\leq|\mu|(\mathbb{C})\epsilon.$
On the other hand,
$\displaystyle\|\sum_{k\in S_{2}}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]k_{a_{k}^{(N)}}\otimes(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\|_{\mathcal{N}}$
$\displaystyle\leq 2|\mu|(\mathbb{C}\setminus\overline{B(0,R)})$
$\displaystyle\leq 2\epsilon.$
Therefore, we have $\|I_{2}\|_{\mathcal{N}}\leq(2+|\mu|{(\mathbb{C}}))\epsilon$. Similarly ,we have $\|I_{1}\|_{\mathcal{N}}\leq C\epsilon$.
Then for any $N^{\prime}>N$, we have
$\|T_{r_{N}}-T_{r_{N}^{\prime}}\|_{\mathcal{N}}\leq\|I_{1}\|_{\mathcal{N}}+\|I_{2}\|_{\mathcal{N}}\leq C\epsilon.$
Thus, $T_{r_{n}}$ is a Cauchy sequence in $\mathcal{N}({f_{\alpha}^{\infty},F_{\alpha}^{1}})$. Then there is a $T\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$, such that $T_{r_{n}}\rightarrow T$ with $\|\cdot\|_{\mathcal{N}}$. Therefore, $T_{r_{n}}\rightarrow T$ with $\|\cdot\|_{op}$ (operator norm), since $\|A\|_{op}\leq\|A\|_{\mathcal{N}}$ for all $A\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$. Then it implies that $\widetilde{T_{r_{n}}}\rightarrow{\widetilde{T}}$ pointwise.
Recall that $\widetilde{\mu}(z)=\frac{\alpha}{\pi}\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}d\mu(w)$, we have
$\displaystyle|\widetilde{T_{r}}(z)-\frac{\pi}{\alpha}\widetilde{\mu}(z)|=|\sum_{k=1}^{\infty}\mu(S_{r}+a_{k})e^{-\alpha|z-a_{k}|^{2}}-\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}d\mu(w)|$
$\displaystyle=|\sum_{k=1}^{\infty}\int_{S_{r}+a_{k}}e^{-\alpha|z-w|^{2}}-e^{-\alpha|z-a_{k}|^{2}}d\mu(w)|$
$\displaystyle\leq\sum_{k=1}^{\infty}\int_{S_{r}+a_{k}}|e^{-\alpha|z-w|^{2}}-e^{-\alpha|z-a_{k}|^{2}}|d|\mu|(w).$
Since $e^{-\alpha|z|^{2}}$ is uniformly continuous on $\mathbb{C}$, for any $\epsilon>0$, there exists a $r_{0}>0$, if $r<r_{0}$, we have
$|e^{-\alpha|z-w|^{2}}-e^{-\alpha|z-a_{k}|^{2}}|<\frac{\epsilon}{|\mu|(\mathbb{C})}$
for every $w\in S_{r}+a_{k}$. Hence, if $r<r_{0}$, we have
$\displaystyle|\widetilde{T_{r}}(z)-\frac{\pi}{\alpha}\widetilde{\mu}(z)|<\epsilon.$
This implies $lim_{r\rightarrow 0}\widetilde{T_{r}}(z)=\frac{\pi}{\alpha}\widetilde{\mu}(z)$, hence $\widetilde{T}(z)=\frac{\pi}{\alpha}\widetilde{\mu}(z)$.
Set $F(z,w)=T_{\mu}K_{z}(w)$ and $G(z,w)=TK_{z}(w)$. These functions are analytic with respect to $w$, and are anti-holomorphic with respect to $z$. Furthermore, we have $F(z,z)=\frac{\alpha}{\pi}G(z,z)=e^{\alpha|z|^{2}}\widetilde{\mu}(z)$. A standard polarization theorem in several complex variables (see [14]) yields $F(z,w)=\frac{\alpha}{\pi}G(z,w)$ for all $z,w\in\mathbb{C}$. Then we have $T_{\mu}K_{z}=\frac{\alpha}{\pi}TK_{z}$ for all $z\in\mathbb{C}$. By the density of the finite linear span of kernel functions in $f_{\alpha}^{\infty}$, we have $T_{\mu}=\frac{\alpha}{\pi}T\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$.
The above proof shows that $\frac{\alpha}{\pi}T_{r_{n}}\stackrel{{\scriptstyle\|\cdot\|_{\mathcal{N}}}}{{\longrightarrow}}T_{\mu}$. Then we have
$\|T_{\mu}\|_{\mathcal{N}}=\frac{\alpha}{\pi}\lim\limits_{n\rightarrow\infty}\|T_{r_{n}}\|_{\mathcal{N}}\leq\frac{\alpha}{\pi}|\mu|(\mathbb{C}).$
Recall that $F_{\alpha}^{1}\subset F_{\alpha}^{p}$ for any $p\geq 1$, and the inclusion is continuous. Therefore, we obtain the following corollary.
Corollary 3.6.
Let $\mu$ be a complex Borel measure on $\mathbb{C}$, and $|\mu|(\mathbb{C})<\infty$. Then $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$, where $1\leq p<\infty,1\leq q\leq\infty$. Moreover
$\|T_{\mu}\|_{\mathcal{N}}\lesssim|\mu|(\mathbb{C}).$
By the hypothesis and Theorem 3.5, we have $T_{\mu}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$. Then we can find $(f_{j})$, $(g_{j})\subset F_{\alpha}^{1}$ such that $T_{\mu}=\sum_{j=1}^{\infty}f_{j}\otimes g_{j}$, and $\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,1}\|g_{j}\|_{\alpha,1}<\infty$. Since $p^{\prime},q>1$, the embedding from $F_{\alpha}^{1}$ to $F_{\alpha}^{p^{\prime}}$ and $F_{\alpha}^{q}$ is continuous. Hence
$\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,p^{\prime}}\|g_{j}\|_{\alpha,q}\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,1}\|g_{j}\|_{\alpha,1}<\infty.$
Then we have $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$, and $\|T_{\mu}\|_{\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})}\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,1}\|g_{j}\|_{\alpha,1}$. Take the infimum over all possible representations of $T_{\mu}$ on the right-hand side of the inequality, we obtain $\|T_{\mu}\|_{\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})}\lesssim\|T_{\mu}\|_{\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})}\lesssim|\mu|(\mathbb{C})$. ∎
In 2020, Fulsche proved that $T_{f}\in\mathcal{N}(F_{\alpha}^{p})$ if $f\in L^{1}(\mathbb{C})$ (see [9, Lemma 2.11]), thus the above corollary can be viewed as a generalization of this result. Moreover, if $\mu$ is a positive measure, the below theorem shows that the condition $\mu(\mathbb{C})<\infty$ is also sufficient for $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ when $q\leq p$.
Theorem 3.7.
Let $1\leq q\leq p<\infty,1\leq s\leq\infty$, and let $\mu\geq 0$ satisfying (1.2), then the following statements are equivalent
  1. (1)
    $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$;
  2. (2)
    $T_{\mu}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{s})$;
  3. (3)
    $\mu(\mathbb{C})<\infty$.
Moreover,
$\|T_{\mu}\|_{\mathcal{N}}\simeq\mu(\mathbb{C}).$ (3.1)
By a direct calculation, we see that $\widetilde{T_{\mu}}(z)=\widetilde{\mu}(z)$. Since $\mu\geq 0$, we can use Fubini’s Theorem to obtain
$\displaystyle\|\widetilde{\mu}\|_{L^{1}}$ $\displaystyle=\frac{\alpha}{\pi}\int_{\mathbb{C}}\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}d\mu(w)dA(z)$
$\displaystyle=\frac{\alpha}{\pi}\int_{\mathbb{C}}\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}dA(z)d\mu(w)$
$\displaystyle=\mu(\mathbb{C}).$
$(1)\Rightarrow(3)$: If $1\leq q\leq p<\infty$, by Theorem 3.2 (3), we have $\widetilde{T_{\mu}}(z)=\widetilde{\mu}(z)\in L^{1}(\mathbb{C},dA)$, which implies $\mu(\mathbb{C})<\infty$. $(3)\Rightarrow(1)$ follows from Corollary 3.6. The proof of $(2)\Leftrightarrow(3)$ is similar.
Estimate (3.1) follows from the equation $\|\widetilde{\mu}\|_{L^{1}}=\mu(\mathbb{C})$, Theorem 3.2 and Theorem 3.5. ∎
Now, we present an interesting result. This conclusion is not immediately obvious, but it follows directly from our previous results.
Corollary 3.8.
Let $f\in L^{1}(\mathbb{C})$, then $\widetilde{f}\in\cap_{r\geq 1}L^{r}(\mathbb{C})$. Moreover, for any $r\geq 1$, we have $\|\widetilde{f}\|_{L^{r}}\lesssim\|f\|_{L^{1}}$.
It follows from Theorem 3.5, we have $T_{f}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$, then by Theorem 3.2, we have $\widetilde{f}=\widetilde{T_{f}}\in L^{r}(\mathbb{C})$ for all $\frac{1}{2}\leq r\leq\infty$. Moreover, applying Theorem 3.2 and Theorem 3.5, we obtain
$\|\widetilde{f}\|_{L^{r}}\lesssim\|T_{f}\|_{\mathcal{N}}\lesssim\|f\|_{L^{1}}.$
Let $\overline{F}_{\alpha}^{p}=\{\overline{f}:f\in F_{\alpha}^{p}\}$, suppose $\mu$ is a complex Borel measure and satisfies condition (1.1), we can define small Hankel operator on $K$ by
$h_{\mu}(f)(z)=\frac{\alpha}{\pi}\int_{\mathbb{C}}K_{\alpha}(w,z)f(w)\mathrm{e}^{-\alpha|w|^{2}}\,d\mu(w),\quad z\in\mathbb{C}.$
Theorem 3.9.
Let $\mu$ be a complex Borel measure on $\mathbb{C}$, and $|\mu|(\mathbb{C})<\infty$. Then $h_{\mu}\in\mathcal{N}(F_{\alpha}^{p},\overline{F}_{\alpha}^{q})(\mathcal{N}(f_{\alpha}^{\infty},\overline{F}_{\alpha}^{q}))$, where $1\leq p<\infty,1\leq q\leq\infty$. Moreover, we have
$\|h_{\mu}\|_{\mathcal{N}}\lesssim|\mu|(\mathbb{C}).$
Following the same method as in the proof of Theorem 3.5, we define
$h_{r}=\sum_{k=1}^{\infty}\mu(S_{r}+a_{k})\,k_{a_{k}}\otimes\overline{k_{a_{k}}}.$
Taking $r_{n}=\frac{1}{2^{n}}$, we have $h_{r_{n}}\to h_{\mu}$ in the nuclear norm $\|\cdot\|_{\mathcal{N}}$, which establishes the desired result.
Definition 3.10.
Let $X$ and $Y$ be Banach spaces. A linear operator $T\colon X\to Y$ is called absolutely summing provided there exists a constant $C>0$ such that for any finite family $\{x_{j}\}_{j=1}^{n}\subset X$,
$\sum_{j=1}^{n}\|Tx_{j}\|\;\leq\;C\sup_{\|x^{\prime}\|\leq 1}\sum_{j=1}^{n}|x^{\prime}(x_{j})|\;=\;C\sup_{|\varepsilon_{j}|=1}\Big\|\sum_{j=1}^{n}\varepsilon_{j}x_{j}\Big\|.$
Every nuclear operator is absolutely summing. In the special cases $X=c_{0}$ or $X=\ell_{\infty}$, the two classes coincide [8, Lemma 2.1].
Corollary 3.11.
Let $1\leq s\leq\infty$, and let $\mu\geq 0$ satisfying (1.2), then $T_{\mu}:f_{\alpha}^{\infty}\rightarrow F_{\alpha}^{s}$ is absolutely summing if and only if $\mu(\mathbb{C})<\infty$.
It is known that $f_{\alpha}^{\infty}$ is isomorphic to $c_{0}$(see [15]). Then the conclusion follows from Theorem 3.7. ∎
Remark 3.12.
Suppose $1\leq p<q,$ and $\mu\geq 0$. If $\mu(\mathbb{C})<\infty$, by Corollary 3.6, we have $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$. However, we cannot use Theorem 3.2 to prove the equivalence between $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ and $\mu(\mathbb{C})<\infty$, since in this case $(\frac{1}{p^{\prime}}+\frac{1}{q})^{-1}>1$. In fact, we will show that there exists $T\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ such that $\widetilde{T}(z)\notin L^{1}$.
Theorem 3.13.
Let $1\leq p<\infty$, and let $g$ is a entire function. Then
$\int_{\mathbb{C}}|f(z)\overline{g(z)}|e^{-\alpha|z|^{2}}dA(z)<\infty$
for all $f\in F_{\alpha}^{p}$ if and only if $g\in F_{\alpha}^{p^{\prime}}$.
$\Rightarrow:$ For any $n\in\mathbb{N}^{+}$, define the truncated function
$g_{n}(z)=\chi_{\{|g(z)|\leq n\}}\cdot g(z),$
so that
$\lim_{n\rightarrow\infty}g_{n}(z)=g(z)$
for all $z\in\mathbb{C}$. For any measurable function $h$, define $\varphi_{h}:F_{\alpha}^{p}\rightarrow\mathbb{C}$ by
$\varphi_{h}(f)=\int_{\mathbb{C}}f(z)\overline{h(z)}e^{-\alpha|z|^{2}}dA(z),\quad f\in F_{\alpha}^{p}.$
Then we have $\varphi_{g_{n}}\in(F_{\alpha}^{p})^{\prime}$. Fix $f\in F_{\alpha}^{p}$, since $|g_{n}(z)|\leq|g(z)|$, we can use dominated convergence theorem to obtain
$\lim_{n\rightarrow\infty}\varphi_{g_{n}}(f)=\varphi_{g}(f).$
Hence $\sup_{n}|\varphi_{g_{n}}(f)|<\infty$. The Principle of Uniform Boundedness yields
$\sup_{n}\|\varphi_{g_{n}}\|=C<\infty.$
Then
$|\varphi_{g}(f)|=\lim\limits_{n\rightarrow\infty}|\varphi_{g_{n}}(f)|\leq C\|f\|$
for all $f\in F_{\alpha}^{p}$. It implies that $\varphi_{g}\in(F_{\alpha}^{p})^{\prime}$. Since we can identify the dual space $(F_{\alpha}^{p})^{\prime}$ with $F_{\alpha}^{p^{\prime}}$ through the pairing $\langle\cdot,\cdot\rangle_{F_{\alpha}^{2}}$, we can find $g^{\prime}\in F_{\alpha}^{p^{\prime}}$ such that
$\varphi_{g-g^{\prime}}=0.$
Suppose $(g-g^{\prime})(z)=\sum\limits_{k=0}^{\infty}a_{k}z^{k}$. For any $n\in\mathbb{N}$ , we have
$\int_{\mathbb{C}}|z^{n}||\overline{(g-g^{\prime})(z)}|e^{-\alpha|z|^{2}}dA(z)<\infty.$
The dominated convergence theorem then gives
$\displaystyle\varphi_{g-g^{\prime}}(z^{n})$ $\displaystyle=\int_{\mathbb{C}}z^{n}\overline{(g-g^{\prime})(z)}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=lim_{R\rightarrow{\infty}}\int_{|z|<R}z^{n}\overline{(g-g^{\prime})(z)}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=lim_{R\rightarrow{\infty}}\overline{a_{n}}\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z).$
$\displaystyle=\overline{a_{n}}\int_{\mathbb{C}}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=0.$
It implies that $a_{n}=0$ for all $n\in\mathbb{N}$. Then we have $g=g^{\prime}\in F_{\alpha}^{p^{\prime}}$.
The other direction follows from Hölder’s inequality. ∎
Example 3.14.
It is well-known that for $1\leq p<q\leq\infty$, we have the (proper) inclusion $F_{\alpha}^{p}\subset F_{\alpha}^{q}$. Theorem 3.13 tells us: We can find $f\in F_{\alpha}^{p^{\prime}}$ and $g\in F_{\alpha}^{q}$ such that
$\int_{\mathbb{C}}|f(z)\overline{g(z)}|e^{-\alpha|z|^{2}}dA(z)=\infty.$
In other words, if we define $T=f\otimes g\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$, then we have $\|\widetilde{T}\|_{L^{1}}=\infty.$ Subsequently, we give an explicit construction of $f$ and $g$.
If $1<p<q<\infty$, we have $1<q^{\prime}<p^{\prime}<\infty$. Let $a=\frac{1}{q}-\frac{1}{p}=\frac{1}{p^{\prime}}-\frac{1}{q^{\prime}}$ and define $\epsilon_{n}=n^{\frac{a}{2}}$. A sequence of indices $n_{1}<n_{2}<\cdots$ can be chosen so that $3^{-k}\leq\epsilon_{n_{k}}\leq 2^{-k}$ for all $k\in\mathbb{N}$. Fix a constant $b$ with $\sqrt{3}<b<2$. We consider the function defined by $f(z)=\sum a_{n}z^{n}$, where its coefficients are given by
$a_{n}=\begin{cases}b^{k}\sqrt{\frac{\alpha^{n_{k}}}{(n_{k})!}}\,n_{k}^{\frac{1}{4}-\frac{1}{2p^{\prime}}}\,\epsilon_{n_{k}},&\text{if }n=n_{k}\text{ for some }k,\\ 0,&\text{otherwise}.\end{cases}$
The membership of $f$ in $F_{\alpha}^{p^{\prime}}$ is verified by computing two sums. For $1<p^{\prime}\leqslant 2$,
$\sum_{n=0}^{\infty}|a_{n}|^{p^{\prime}}\left(\frac{n!}{\alpha^{n}}\right)^{\frac{p^{\prime}}{2}}n^{-\frac{p^{\prime}}{4}+\frac{1}{2}}=\sum_{k=1}^{\infty}b^{p^{\prime}k}\epsilon_{n_{k}}^{p^{\prime}}\leq\sum_{k=1}^{\infty}\left(\frac{b^{p^{\prime}}}{2^{p^{\prime}}}\right)^{k}<\infty.$
If $2\leqslant p^{\prime}<\infty$, then
$\sum_{n=1}^{\infty}\left|a_{n}\right|^{p}\left(\frac{n!}{\alpha^{n}}\right)^{\frac{p}{2}}n^{\frac{p}{4}-\frac{1}{2}}=\sum_{k=1}^{\infty}b^{pk}n_{k}^{\frac{p}{4}-\frac{1}{2}+\frac{p}{4}-\frac{p}{2p^{\prime}}}\epsilon_{n_{k}}^{p}\leq\sum_{k=1}^{\infty}\left(\frac{b^{p}}{2^{p}}\right)^{k}<\infty.$
The convergence of these series, together with Theorems 1 and 4 in [18], implies $f\in F_{\alpha}^{p^{\prime}}.$
Note that the coefficients satisfy
$a_{n}=\begin{cases}b^{k}\sqrt{\frac{\alpha^{n_{k}}}{(n_{k})!}}\,n_{k}^{\frac{1}{4}-\frac{1}{2q^{\prime}}}\,,&\text{if }n=n_{k},\\ 0,&\text{otherwise}.\end{cases}$
Consequently,
$a_{n}\neq o\left(\sqrt{\frac{\alpha^{n}}{n!}}n^{\frac{1}{4}-\frac{1}{2q^{\prime}}}\right)\quad(n\rightarrow\infty),$
which, by Corollary 5 of [18], means $f\in F_{\alpha}^{p^{\prime}}\setminus F_{\alpha}^{q^{\prime}}$.
A function $g$ is constructed analogously. Let
$b_{m}=\begin{cases}b^{k}\sqrt{\frac{\alpha^{n_{k}}}{(n_{k})!}}\,n_{k}^{\frac{1}{4}-\frac{1}{2q}}\,\epsilon_{n_{k}},&\text{if }m=n_{k},\\ 0,&\text{otherwise},\end{cases}$
and define $g(z)=\sum b_{m}z^{m}$. It then follows that $g\in F_{\alpha}^{q}\setminus F_{\alpha}^{p}$.
If $\int_{\mathbb{C}}|f(z){g(z)}|e^{-\alpha|z|^{2}}dA(z)<\infty$, Applying the dominated convergence theorem, we obtain
$\displaystyle\int_{\mathbb{C}}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}dA(z)$ $\displaystyle=lim_{R\rightarrow{\infty}}\int_{|z|<R}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=\lim_{R\rightarrow{\infty}}\sum_{n=0}^{\infty}\sum_{m=0}^{\infty}a_{n}\overline{b_{m}}\int_{|z|<R}z^{n}\overline{z}^{m}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=\lim_{R\rightarrow{\infty}}\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z).$
From the construction above, we have $a_{n}\overline{b_{n}}\geq 0$ and $\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$ is monotonically increasing in $R$. The monotone convergence theorem yields:
$\displaystyle\int_{\mathbb{C}}f(z)\overline{g(z)}e^{\alpha|z|^{2}}dA(z)$ $\displaystyle=lim_{R\rightarrow{\infty}}\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}\int_{\mathbb{C}}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=\frac{\pi}{\alpha}\sum_{k=1}^{\infty}b^{2k}\epsilon_{n_{k}}$
$\displaystyle\geq\frac{\pi}{\alpha}\sum_{k=1}^{\infty}(\frac{b^{2}}{3})^{k}$
$\displaystyle=\infty.$
But $\int_{\mathbb{C}}f(z)\overline{g(z)}\mathrm{e}^{-\alpha|z|^{2}}dA(z)\leq\int_{\mathbb{C}}|f(z){g(z)}|\mathrm{e}^{-\alpha|z|^{2}}dA(z)$. We obtain a contradiction.
However, for $1\leq p<q\leq\infty$ and $\mu\geq 0$, it remains unknown whether $\mu(\mathbb{C})<\infty$ is necessary for $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$. Finally, we conclude this section with an open problem.
Question 3.15.
For $1\leq p<q\leq\infty$ and $\mu\geq 0$, what is the necessary and sufficient condition for $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$?
4. Trace formula and density theorem
In this section, for an operator $T\in\mathcal{N}(F_{\alpha}^{p})$, $(1<p<\infty)$, we will provide some formulas for calculating its trace. Finally, we will establish a density theorem.
Lemma 4.1.
Let $1<p<\infty$ and $T\in\mathcal{N}(F^{p}_{\alpha})$, then we have:
$Tr(T)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{T}(z)dA(z).$
Let $T=\sum_{j=1}^{\infty}f_{j}\otimes g_{j}$, where $\left(g_{j}\right)\subset F_{\alpha}^{p},\left(f_{j}\right)\subset F_{\alpha}^{p^{\prime}}$ be a nuclear representation of $T$. Then we have $Tr(T)=\sum_{j=1}^{\infty}\langle g_{j},f_{j}\rangle=\sum_{j=1}^{\infty}\langle g_{j},f_{j}\rangle_{F_{\alpha}^{2}}$. And the trace is independent of the representation choice. Then we have:
$\displaystyle\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{T}(z)dA(z)$ $\displaystyle=\frac{\alpha}{\pi}\int_{\mathbb{C}}\langle Tk_{z},k_{z}\rangle_{F_{\alpha}^{2}}dA(z)$
$\displaystyle=\frac{\alpha}{\pi}\int_{\mathbb{C}}\sum_{j=1}^{\infty}\langle K_{z},f_{j}\rangle_{F_{\alpha}^{2}}\langle g_{j},K_{z}\rangle_{F_{\alpha}^{2}}e^{-\alpha|z|^{2}}dA(z)$
$\displaystyle=\sum_{j=1}^{\infty}\int_{\mathbb{C}}\overline{f_{j}(z)}g_{j}(z)d\lambda_{\alpha}(z)$
$\displaystyle=\sum_{j=1}^{\infty}\langle g_{j},f_{j}\rangle_{F_{\alpha}^{2}}$
$\displaystyle=Tr(T).$
Proposition 4.2.
Suppose $\mu\geq 0$ and $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p})$, where $1<p<\infty$, then $Tr(T_{\mu})=\frac{\alpha}{\pi}\mu(\mathbb{C})$.
Let $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p})$, then we have $\mu(\mathbb{C})<\infty$ . By Fubini’s theorem, we obtain:
$Tr(T_{\mu})$ $=\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{\mu}(z)dA(z)$
$=(\frac{\alpha}{\pi})^{2}\int_{\mathbb{C}}dA(z)\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}d\mu(w)$
$=(\frac{\alpha}{\pi})^{2}\int_{\mathbb{C}}d\mu(w)\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}dA(z)$
$=\frac{\alpha}{\pi}\mu(\mathbb{C}).$
We now prove a useful result. The technical hypotheses in its statement are needed to justify the application of Fubini’s theorem. Notice that the case $p=q=2$ of the following theorem appears in [3, Theorem 6], but we adopt a different approach here.
Theorem 4.3.
Suppose $\varphi$ is Lebesgue measurable on $\mathbb{C}$ and $S\in\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p})$, where $1<p,q<\infty$. If
  1. (1)
    $\varphi K_{z}\in L_{\alpha}^{p^{\prime}}$ for every $z\in\mathbb{C}$,
  2. (2)
    $T_{\varphi}\in\mathcal{L}(F_{\alpha}^{p},F_{\alpha}^{q})$,
  3. (3)
    $T_{\varphi}S\in\mathcal{N}(F_{\alpha}^{q})$,
  4. (4)
    $\int_{\mathbb{C}}\int_{\mathbb{C}}|\varphi(w)||K(z,w)||SK_{z}(w)|d\lambda_{\alpha}(z)d\lambda_{\alpha}(w)<\infty$.
Then we have
$Tr(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(z)\widetilde{S}(z)dA(z).$
Let $f\in F_{\alpha}^{p}$, we can find a sequence $f_{n}\in K$ such that $f_{n}$ converges to $f$ with $\|\|_{\alpha,p}$. The set $K$ is defined as (1.3). By assumption (2), we have $T_{\varphi}f_{n}\stackrel{{\scriptstyle\|\cdot\|_{\alpha,q}}}{{\longrightarrow}}T_{\varphi}f$, hence $T_{\varphi}f_{n}(z)\rightarrow T_{\varphi}f(z)$ for every $z\in\mathbb{C}$.
Since $f_{n}\in K$, we have
$T_{\varphi}f_{n}(z)=\int_{\mathbb{C}}f_{n}(w)\varphi(w)K_{\alpha}(z,w)d\lambda_{\alpha}(w)$
for all $n\in\mathbb{N}$ and $z\in\mathbb{C}$. Then fix $z\in\mathbb{C}$, by assumption (1), we have
$\displaystyle|\int_{\mathbb{C}}(f_{n}(w)-f(w))\varphi(w)K(z,w)d\lambda_{\alpha}(w)|$
$\displaystyle\leq\frac{\alpha}{\pi}\int_{\mathbb{C}}|(f_{n}(w)-f(w))e^{-\frac{\alpha}{2}|z|^{2}}||\varphi(w)K(z,w)e^{-\frac{\alpha}{2}|z|^{2}}|dA(w)$
$\displaystyle\lesssim\|f_{n}-f\|_{\alpha,p}\|\varphi K_{z}\|_{\alpha,p^{\prime}}\rightarrow 0(n\rightarrow\infty).$
Hence
$T_{\varphi}f(z)=\int_{\mathbb{C}}f(w)\varphi(w)K(z,w)d\lambda_{\alpha}(w)$
for every $f\in F_{\alpha}^{p}$. Then we can use Fubini’s Theorem to obtain
$Tr(T_{\varphi}S)$ $=\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{T_{\varphi}S}(z)dA(z)$ (4.1)
$=\int_{\mathbb{C}}\langle T_{\varphi}SK_{z},K_{z}\rangle_{F_{\alpha}^{2}}d\lambda_{\alpha}(z)$
$=\int_{\mathbb{C}}T_{\varphi}SK_{z}(z)d\lambda_{\alpha}(z)$
$=\int_{\mathbb{C}}d\lambda_{\alpha}(z)\int_{\mathbb{C}}\varphi(u)K_{u}(z)SK_{z}(u)d\lambda_{\alpha}(u)$
$=\int_{\mathbb{C}}\varphi(u)d\lambda_{\alpha}(u)\int_{\mathbb{C}}K_{u}(z)SK_{z}(u)d\lambda_{\alpha}(z).$
Let $S^{\prime}\in\mathcal{L}(F_{\alpha}^{p^{\prime}},F_{\alpha}^{q^{\prime}})$ be the adjoint operator of $S$ under the pairing $\langle\cdot,\cdot\rangle_{F_{\alpha}^{2}}$. Then we have
$\langle Sf,g\rangle_{F_{\alpha}^{2}}=\langle f,S^{\prime}g\rangle_{F_{\alpha}^{2}}$
for all $f\in F_{\alpha}^{q},g\in F_{\alpha}^{p^{\prime}}$ . Hence:
$\displaystyle\int_{\mathbb{C}}K_{u}(z)SK_{z}(u)d\lambda_{\alpha}(z)$ $\displaystyle=\int_{\mathbb{C}}K_{u}(z)\langle SK_{z},K_{u}\rangle_{F_{\alpha}^{2}}d\lambda_{\alpha}(z)$
$\displaystyle=\int_{\mathbb{C}}K_{u}(z)\langle K_{z},S^{\prime}K_{u}\rangle_{F_{\alpha}^{2}}d\lambda_{\alpha}(z)$
$\displaystyle=\overline{\int_{\mathbb{C}}K_{z}(u)S^{\prime}K_{u}(z)d\lambda_{\alpha}(z)}$
$\displaystyle=\overline{S^{\prime}K_{u}(u)}$
$\displaystyle=SK_{u}(u)$
$\displaystyle=e^{\alpha|u|^{2}}\widetilde{S}(u).$
It follows from (4.1), we have:
$Tr_{F_{\alpha}^{q}}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(u)\widetilde{S}(u)dA(u).$
∎ We can now prove
Lemma 4.4.
Let $1<p,q<\infty$, suppose $\varphi$ is a bounded, compactly supported function on $\mathbb{C}$. Then $T_{\varphi}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$. Furthermore, for any $S\in\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p})$ we have:
$Tr_{F_{\alpha}^{q}}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(z)\widetilde{S}(z)dA(z).$ (4.2)
$T_{\varphi}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ follows from Corollary 3.6. Applying the pointwise estimates on the Fock space, we have
$|SK_{z}(u)|\leq\|SK_{z}\|_{\alpha,p}e^{\frac{\alpha}{2}|u|^{2}}\leq\|S\|e^{\frac{\alpha}{2}|z|^{2}}e^{\frac{\alpha}{2}|u|^{2}}.$
Hence
$\displaystyle\int_{\mathbb{C}}|\varphi(u)|d\lambda_{\alpha}(u)\int_{\mathbb{C}}|K_{u}(z)||SK_{z}(u)|d\lambda_{\alpha}(z)$
$\displaystyle\lesssim\int_{\mathbb{C}}|\varphi(u)|e^{\frac{\alpha}{2}|u|^{2}}d\lambda_{\alpha}(u)\int_{\mathbb{C}}|K_{u}(z)|e^{-\frac{\alpha}{2}|z|^{2}}dA(z)$
$\displaystyle\simeq\int_{\mathbb{C}}|\varphi(u)|dA(u)$
$\displaystyle<\infty.$
Then the conclusion follows form Theorem 4.3.
The theorem below extends a result in [3, Theorem 9]. The cited theorem asserts that the collection of all Toeplitz operators $T_{\varphi}$ with compactly supported continuous symbols is both trace-norm dense in the trace ideal of $F_{\alpha}^{2}$ and operator-norm dense in $\mathcal{K}(F_{\alpha}^{2})$.
Theorem 4.5.
Let $\mathcal{C}$ denote the set of all Toeplitz operators $T_{\varphi}$ with continuous, compactly supported symbols $\varphi$ on $\mathbb{C}$, and let $1<p,q<\infty$. Then:
  1. (1)
    $\mathcal{C}$ is dense in $\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ with respect to the nuclear norm.
  2. (2)
    $\mathcal{C}$ is norm-dense in $\mathcal{K}(F_{\alpha}^{p},F_{\alpha}^{q})$.
We first prove (2). Assume, for contradiction, that $\mathcal{C}$ is not norm-dense in $\mathcal{K}(F_{\alpha}^{p},F_{\alpha}^{q})$. An application of the Hahn–Banach theorem combined with Corollary 2.5 implies the existence of a nonzero operator $S\in\mathcal{N}(F_{\alpha}^{q},F_{\alpha}^{p})$ satisfying
$\operatorname{Tr}(T_{\varphi}S)=0\qquad\text{for all }T_{\varphi}\in\mathcal{C}.$
By Lemma 4.4, this trace can be written as
$\operatorname{Tr}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(z)\,\widetilde{S}(z)\,dA(z).$
Since the equality holds for every continuous, compactly supported $\varphi$, it follows that $\widetilde{S}(z)=0$. Consequently, $S=0$, which contradicts the fact that $S$ is nonzero. This establishes statement (2).
The proof of (1) follows a similar argument. ∎
Funding
Y. Lu was supported by NNSFC (Grant No.12031002), C. Zu was supported by NNSFC (Grant No.12401151), and the Postdoctoral Researcher Foundation of China (Grant No.GZB20240100).
Data availability statement
No data were collected, generated or consulted in connection with this research.
Declarations
Conflict of interest
The authors have no conflict of interest.

### v2

@@
 Nuclear Toeplitz Operators between Fock Spaces
-Abstract.
-We study Toeplitz operators with measure-valued symbols acting between Fock spaces.
-Given $1\leq p,q\leq\infty$ and a Borel measure $\mu$ on $\mathbb{C}$, we investigate when the associated Toeplitz operator
-$T_{\mu}:F^{p}_{\alpha}\to F^{q}_{\alpha}$
-belongs to the nuclear class.
-For positive measures $\mu$ and in the range $1\leq q\leq p\leq\infty$, we obtain necessary and sufficient conditions for the nuclearity of $T_{\mu}$ in terms of the Berezin transform of $\mu$.
-As a consequence, nuclearity in this setting exhibits a rigidity property: if $T_{\mu}$ is nuclear from $F^{p}_{\alpha}$ to $F^{q}_{\alpha}$ for some $q\leq p$, then it is nuclear for all such $q$.
-In the case $p<q$, we show that the situation is more delicate.
-We provide separate necessary and sufficient conditions for nuclearity, indicating that the Berezin transform alone does not yield a complete characterization.
-The proofs rely on tools from Banach space operator theory combined with kernel estimates on Fock spaces.
-Our results extend naturally to Fock spaces on $\mathbb{C}^{n}$.
-Key words and phrases:
+Abstract
+We characterize the nuclearity of Toeplitz operators $T_{\mu}:F_{\alpha}^{p}\to F_{\alpha}^{q}$ with Borel measure symbols for $1\leq p,q\leq\infty$.
+For positive measures $\mu$ and $q\leq p$, we provide necessary and sufficient conditions in terms of the Berezin transform and establish a rigidity property for nuclearity across this range.
+In the case $p<q$, we obtain separate necessary and sufficient conditions, indicating that the Berezin transform alone is insufficient for a complete characterization.
+Our results extend to Fock spaces on $\mathbb{C}^{n}$.
+keywords:
 Nuclear operator, Toeplitz operator, Fock space
-1991 Mathematics Subject Classification:
-Primary 47B10 Secondary 47B35
-*Corresponding author.
-1. Introduction
-Let $\alpha>0$ and $p>0$, the notation $L_{\alpha}^{p}$ refers to the space of all Lebesgue measurable functions $f$ on $\mathbb{C}$ for which $f(z)\mathrm{e}^{-\alpha|z|^{2}/2}\in L^{p}(\mathbb{C},\mathrm{d}A)$, where $\mathrm{d}A$ is the Lebesgue area measure on $\mathbb{C}$.
-For $f\in L_{\alpha}^{p}$, its norm is given by
-$\|f\|_{\alpha,p}^{p}=\frac{p\alpha}{2\pi}\int_{\mathbb{C}}\left|f(z)\mathrm{e}^{-\frac{\alpha}{2}|z|^{2}}\right|^{p}\,dA(z).$
-Correspondingly, for $\alpha>0$ and $p=\infty$, we denote by $L_{\alpha}^{\infty}$ the space of measurable functions $f$ such that
-$\|f\|_{\infty,\alpha}=\operatorname{ess\,sup}\left\{|f(z)|\mathrm{e}^{-\alpha|z|^{2}/2}:z\in\mathbb{C}\right\}<\infty.$
+articletype: RESEARCH ARTICLE
+{Frontmatter}
+1 Introduction
+Toeplitz operators on Fock spaces constitute a fundamental class of non-selfadjoint operators in complex analysis and operator theory and have attracted a great deal of interest.
+The theory of Toeplitz operators on Fock spaces has been extensively developed in the Hilbert space setting.
+For the classical Fock space $F_{\alpha}^{2}$, Toeplitz operators with bounded or measure-valued symbols have been studied from various perspectives, including boundedness, compactness, and Schatten class membership, see references [MR2610379, MR882716, MR1277446, MR4107813].
+In particular, Isralowitz and Zhu [MR2609242] obtained complete characterizations of boundedness, compactness, and Schatten class properties of Toeplitz operators with positive Borel measure symbols.
+Beyond the Hilbert space situation, Hu and Lv [MR2819157, MR3248473] systematically investigated Toeplitz operators acting between different Fock spaces $F_{\alpha}^{p}$ and $F_{\alpha}^{q}$ for $0<p,q<\infty$.
+They characterized bounded and compact Toeplitz operators with positive Borel measure symbols in terms of averaging functions and Berezin-type transforms.
+A notable feature of their results is a rigidity phenomenon specific to Fock spaces: boundedness (or compactness) from $F_{\alpha}^{p}$ to $F_{\alpha}^{q}$ for some $p\leq q$ automatically implies boundedness (or compactness) for all such $q$.
+The groundbreaking work of Hu and Lv has stimulated extensive scholarly investigations into Toeplitz and Hankel operators across diverse weighted Fock spaces [MR3392258, MR3157978, MR2934601].
+In contrast to the well-developed theory of bounded and compact Toeplitz operators, considerably less is known about Toeplitz operators belonging to finer operator ideals on Fock spaces, especially in the non-Hilbertian case.
+Among these ideals, the class of nuclear operators plays a distinguished role in Banach space operator theory.
+The aim of this paper is to study nuclear Toeplitz operators acting between Fock spaces.
+Our first main result is presented as follows:
+Theorem 1.1.
+Let $1\leq q\leq p<\infty,1\leq s\leq\infty$, and let $\mu\geq 0$ satisfy (2.1); then the following statements are equivalent
+  1. (1)
+    $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$;
+  2. (2)
+    $T_{\mu}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{s})$;
+  3. (3)
+    $\mu(\mathbb{C})<\infty$.
+Moreover,
+$\|T_{\mu}\|_{\mathcal{N}}\simeq\mu(\mathbb{C}).$ (1.1)
+Since nuclear and trace-class operators coincide on Hilbert spaces, the case $p=q=2$ in our results recovers Zhu’s characterization of trace‑class Toeplitz operators in [MR2609242].
+Notice that Zhu’s proof relies heavily on Hilbert space structure, including orthogonal decompositions, positivity, and the classical Schatten class theory.
+In contrast, when $p\neq 2$, the Fock space $F^{p}_{\alpha}$ is a Banach space lacking Hilbert geometry.
+The standard techniques used in the Hilbert case do not extend directly.
+Thus, the characterization of nuclear Toeplitz operators between Banach Fock spaces requires a different approach based on operator ideals.
+We emphasize that our method provides a more constructive approach to approximating $T_{\mu}$ with nuclear operators compared to Zhu’s proof.
+Our main results provide a detailed description of nuclear Toeplitz operators in different parameter regimes.
+When $1\leq q\leq p<\infty$ and $\mu\geq 0$, we obtain necessary and sufficient conditions for $T_{\mu}$ to be nuclear in terms of the integrability of the Berezin transform of $\mu$.
+As a consequence, we show that nuclearity in this range exhibits a rigidity property analogous to that observed for boundedness: if $T_{\mu}$ is nuclear from $F_{\alpha}^{p}$ to $F_{\alpha}^{q}$ for some $q\leq p$, then it is nuclear for all such $q$.
+When $p<q$, the situation becomes more subtle.
+In this case, we establish separate necessary and sufficient conditions for nuclearity, showing that the Berezin transform alone does not yield a complete characterization.
+This reflects an inherent asymmetry between the domain and target spaces in the nuclear setting and distinguishes the nuclear theory from previously studied mapping properties of Toeplitz operators on Fock spaces.
+We will discuss this problem in Chapter 3.
+Berger and Coburn [MR1277446] showed that the set of all Toeplitz operators $T_{\varphi}$ with compact support continuous symbol is trace-norm dense in the trace ideal of $F_{\alpha}^{2}$, and is operator-norm dense in $\mathcal{K}(F_{\alpha}^{2})$.
+Our second main result generalizes the aforementioned finding :
+Theorem 1.2.
+Let $C$ denote the set of all Toeplitz operators $T_{\varphi}$ with continuous, compactly supported symbols $\varphi$ on $\mathbb{C}$, and let $1<p,q<\infty$. Then:
+  1. (1)
+    $C$ is dense in $\mathcal{N}(F^{p}_{\alpha},F^{q}_{\alpha})$ with respect to the nuclear norm.
+  2. (2)
+    $C$ is norm-dense in $\mathcal{K}(F^{p}_{\alpha},F^{q}_{\alpha})$.
+The proofs rely on tools from Banach space operator theory, including the duality properties of nuclear operators, approximation and Radon-Nikodym properties of Fock spaces, which we will present in Section 2.
+The paper is organized as follows.
+In Section 2, we recall basic definitions and preliminary results concerning nuclear operators and Fock spaces.
+Section 3 is devoted to the characterization of nuclear Toeplitz operators between Fock spaces and contains our main results.
+In Section 4, we present several trace formulas and establish the density theorem.
+2 Preliminary
+2.1 Fock Space
+Let $\alpha>0$ and $p>0$, the notation $L^{p}_{\alpha}$ refers to the space of all Lebesgue measurable functions $f$ on $\mathbb{C}$ for which $f(z)e^{-\alpha|z|^{2}/2}\in L^{p}(\mathbb{C},\mathrm{d}A)$, where $\mathrm{d}A$ is the Lebesgue area measure on $\mathbb{C}$.
+For $f\in L^{p}_{\alpha}$, its norm is given by
+∥f∥_α,p^p=pα2π∫_C|f(z)e^-α2|z|^2|^pdA(z). Correspondingly, for $\alpha>0$ and $p=\infty$, we denote by $L_{\alpha}^{\infty}$ the space of measurable functions $f$ such that
+∥f∥_∞,α=ess  sup{|f(z)|e^-α|z|^2/2:z∈C}<∞.
 The Fock space $F_{\alpha}^{p}$ is defined as the subspace of $L_{\alpha}^{p}$ consisting of entire functions.
@@
 Let $f_{\alpha}^{\infty}$ be the closed subspace of $F_{\alpha}^{\infty}$ consisting of $f$ satisfying
-$\lim_{|z|\to\infty}f(z)e^{-\frac{\alpha}{2}|z|^{2}}=0.$
-When $p=2$, the space $F_{\alpha}^{2}$ is a Hilbert space equipped with the inner product
-$\langle f,g\rangle_{F_{\alpha}^{2}}=\frac{\alpha}{\pi}\int_{\mathbb{C}}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}\,dA(z).$
+lim_|z|→∞f(z)e^-α2|z|^2=0. When $p=2$ , the space $F_{\alpha}^{2}$ is a Hilbert space equipped with the inner product
+⟨f,g⟩_F_α^2=απ∫_Cf(z)g(z)e^-α|z|^2 dA(z).
 As a reproducing kernel Hilbert space, its reproducing kernel is $K_{\alpha}(z,w)=e^{\alpha z\overline{w}}$.
@@
 For $1\leq p<\infty$ , let $p^{\prime}$ be the conjugate exponent satisfying $\frac{1}{p}+\frac{1}{p^{\prime}}=1$.
-We can identify the dual of $F_{\alpha}^{p}$ with $F_{\alpha}^{p^{\prime}}$, and the dual of $f_{\alpha}^{\infty}$ with $F_{\alpha}^{1}$, via the integral pairing
-$\langle f,g\rangle_{F_{\alpha}^{2}}=\frac{\alpha}{\pi}\int_{\mathbb{C}}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}\,dA(z),$
-and under this identification, the norms on the dual spaces and the Fock spaces are equivalent (see [19, Corollary 2.25, Theorem 2.26]).
+We can identify the dual of $F_{\alpha}^{p}$ with $F_{\alpha}^{p^{\prime}}$ , and the dual of $f_{\alpha}^{\infty}$ with $F_{\alpha}^{1}$ , via the integral pairing ⟨f,g⟩_F_α^2=απ∫_Cf(z)g(z)e^-α|z|^2 dA(z), and under this identification, the norms on the dual spaces and the Fock spaces are equivalent (see [MR2934601, Corollary 2.25, Theorem 2.26]).
 Given $\alpha>0$ , consider the Gaussian measure
-$d\lambda_{\alpha}(z)=\frac{\alpha}{\pi}e^{-\alpha|z|^{2}}\,dA(z).$
-Define the integral operator $P_{\alpha}$ on $L_{\alpha}^{p}$ by
-$P_{\alpha}f(z)=\int_{\mathbb{C}}K_{\alpha}(z,w)f(w)\,d\lambda_{\alpha}(w),\quad f\in L_{\alpha}^{p}.$
-It is known that $P_{\alpha}:L_{\alpha}^{p}\to F_{\alpha}^{p}$ is bounded for all $p$, and it acts as the identity on $F_{\alpha}^{p}$, that is, $P_{\alpha}f=f$ for $f\in F_{\alpha}^{p}$ (see [19, Corollary 2.22]).
+dλ_α(z)=απe^-α|z|^2 dA(z). Define the integral operator $P_{\alpha}$ on $L_{\alpha}^{p}$ by
+P_αf(z)=∫_CK_α(z,w)f(w) dλ_α(w), f∈L_α^p.
+It is known that $P_{\alpha}:L_{\alpha}^{p}\to F_{\alpha}^{p}$ is bounded for all $p$ , and it acts as the identity on $F_{\alpha}^{p}$ , that is, $P_{\alpha}f=f$ for $f\in F_{\alpha}^{p}$ (see [MR2934601, Corollary 2.22]).
 For a function $\varphi\in L^{\infty}(\mathbb{C})$ , the associated Toeplitz operator $T_{\varphi}:F_{\alpha}^{2}\to F_{\alpha}^{2}$ is defined by
-$T_{\varphi}(f)=P_{\alpha}(\varphi f),\quad f\in F_{\alpha}^{2}.$
-More generally, if $\mu$ is a complex Borel measure on $\mathbb{C}$, the Toeplitz operator $T_{\mu}$ is given by
-$T_{\mu}(f)(z)=\frac{\alpha}{\pi}\int_{\mathbb{C}}K_{\alpha}(z,w)f(w)\mathrm{e}^{-\alpha|w|^{2}}\,d\mu(w),\quad z\in\mathbb{C},$
-provided that the integrability condition
-$\int_{\mathbb{C}}|K_{\alpha}(z,w)|\mathrm{e}^{-\alpha|w|^{2}}\,d|\mu|(w)<\infty$ (1.1)
-holds for every $z\in\mathbb{C}$. Due to the exponential nature of the kernel, condition (1.1) is equivalent to
-$\int_{\mathbb{C}}|K_{\alpha}(z,w)|^{2}\mathrm{e}^{-\alpha|w|^{2}}\,d|\mu|(w)<\infty,\quad z\in\mathbb{C}.$ (1.2)
-Let
+T_φ(f)=P_α(φf), f∈F_α^2. More generally, if $\mu$ is a complex Borel measure on $\mathbb{C}$ , the Toeplitz operator $T_{\mu}$ is given by
+T_μ(f)(z)=απ∫_CK_α(z,w)f(w)e^-α|w|^2 dμ(w), z∈C, provided that the integrability condition
+$\int_{\mathbb{C}}|K_{\alpha}(z,w)|e^{-\alpha|w|^{2}}\,d|\mu|(w)<\infty$ (2.1)
+holds for every $z\in\mathbb{C}$. Let
 $K=\operatorname{span}\{K_{z}:z\in\mathbb{C}\}.$ (2.2)
 The set $K$ is dense in $F_{\alpha}^{p}$ for every $1\leq p<\infty$ and also in $f_{\alpha}^{\infty}$.
-Under assumption (1.2), the operator $T_{\mu}$ is well-defined on $K$ by the Cauchy-Schwarz inequality, and hence on a dense subset of each $F_{\alpha}^{p}$ ($1\leq p<\infty$) and $f_{\alpha}^{\infty}$.
+Under assumption (2.1), the operator $T_{\mu}$ is well-defined on $K$.
 For a measure $\mu$ satisfying (2.1), its Berezin transform is defined as
 $\widetilde{\mu}(z)=\frac{\alpha}{\pi}\int_{\mathbb{C}}|k_{z}(w)|^{2}\mathrm{e}^{-\alpha|w|^{2}}\,d\mu(w)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\mathrm{e}^{-\alpha|z-w|^{2}}\,d\mu(w).$ (2.3)
-Toeplitz operators on Fock spaces form an important class of operators connecting complex analysis, operator theory, and the theory of Banach function spaces.
-Owing to the reproducing kernel structure of Fock spaces, Toeplitz operators provide a natural framework for studying how analytic and measure-theoretic properties of symbols are reflected in operator-theoretic behavior.
-The theory of Toeplitz operators on Fock spaces has been extensively developed in the Hilbert space setting.
-For the classical Fock space $F^{2}_{\alpha}$, Toeplitz operators with bounded or measure-valued symbols have been studied from various perspectives, including boundedness, compactness, and Schatten class membership, see references [1, 2, 3, 9].
-In particular, Isralowitz and Zhu [13] obtained complete characterizations of boundedness, compactness, and Schatten class properties of Toeplitz operators with positive Borel measures as symbols.
-Beyond the Hilbertian framework, Hu and Lv [11, 12] systematically investigated Toeplitz operators acting between different Fock spaces $F^{p}_{\alpha}$ and $F^{q}_{\alpha}$ for $1<p,q<\infty$.
-They characterized boundedness and compactness of Toeplitz operators with positive Borel measure symbols in terms of averaging functions and Berezin-type transforms.
-A notable feature of their results is a rigidity phenomenon specific to Fock spaces: boundedness (or compactness) from $F^{p}_{\alpha}$ to $F^{q}_{\alpha}$ for some $p\leq q$ automatically implies boundedness (or compactness) for all such $q$.
-This behavior differs substantially from what is known in the Bergman space setting.
-The groundbreaking work of Hu and Lv has stimulated extensive scholarly investigations into Toeplitz and Hankel operators across diverse weighted Fock spaces [16, 19].
-In contrast to the well-developed theory of bounded and compact Toeplitz operators, considerably less is known about Toeplitz operators belonging to finer operator ideals on Fock spaces, especially in the non-Hilbertian case.
-Among these ideals, the class of nuclear operators plays a distinguished role in Banach space operator theory.
-Nuclearity represents a strong form of compactness and depends sensitively on the geometry of the underlying Banach spaces.
-As a result, techniques based on Hilbert space methods or Schatten class theory are generally insufficient to address nuclearity questions when $p\neq 2$.
-The aim of this paper is to study nuclear Toeplitz operators acting between Fock spaces.
-Given $1\leq p,q\leq\infty$ and a Borel measure $\mu$ on $\mathbb{C}$, we investigate conditions under which the Toeplitz operator
-$T_{\mu}:F^{p}_{\alpha}\longrightarrow F^{q}_{\alpha}$
-belongs to the nuclear class $N(F^{p}_{\alpha},F^{q}_{\alpha})$.
-Our focus is primarily on Toeplitz operators with positive symbols, although several results extend to complex measures under suitable assumptions.
-Our main results provide a detailed description of nuclear Toeplitz operators in different parameter regimes.
-When $1\leq q\leq p\leq\infty$ and $\mu\geq 0$, we obtain necessary and sufficient conditions for $T_{\mu}$ to be nuclear in terms of the integrability of the Berezin transform of $\mu$ (Theorem 3.7).
-As a consequence, we show that nuclearity in this range exhibits a rigidity property analogous to that observed for boundedness: if $T_{\mu}$ is nuclear from $F^{p}_{\alpha}$ to $F^{q}_{\alpha}$ for some $q\leq p$, then it is nuclear for all such $q$.
-When $p<q$, the situation becomes more subtle.
-In this case, we establish separate necessary and sufficient conditions for nuclearity, showing that the Berezin transform alone does not yield a complete characterization.
-This reflects an inherent asymmetry between the domain and target spaces in the nuclear setting and distinguishes the nuclear theory from previously studied mapping properties of Toeplitz operators on Fock spaces.
-The proofs rely on tools from Banach space operator theory, including duality properties of nuclear operators, approximation and Radon–Nikodým properties of Fock spaces, and precise kernel estimates associated with the Berezin transform.
-The arguments presented here extend naturally to Fock spaces on $\mathbb{C}^{n}$, and the main results remain valid in higher dimensions without essential changes.
-The paper is organized as follows.
-In Section 2, we recall basic definitions and preliminary results concerning nuclear operators and Fock spaces.
-Section 3 is devoted to the characterization of nuclear Toeplitz operators between Fock spaces and contains our main results.
-In Section 4, we prove density results showing that Toeplitz operators with compactly supported continuous symbols are dense in the space of nuclear operators between Fock spaces with respect to the nuclear norm.
-As a consequence, we recover and extend known results on trace ideals in the Hilbert space case $p=2$.
-2. Preliminaries
-This section collects known results and establishes basic facts concerning nuclear operators, which will be used in the subsequent sections.
+2.2 Nuclear Operator
 Let $X$ and $Y$ be complex Banach spaces.
@@
 The dual space of a Banach space $X$ is denoted by $X^{\prime}$.
-For $x\in X$ and $x^{\prime}\in X^{\prime}$, we write $\langle x,x^{\prime}\rangle$ for the dual pairing $x^{\prime}(x)$.We say that an operator $T\in\mathcal{L}(X,Y)$ is nuclear provided that it admits a representation of the form
+For $x\in X$ and $x^{\prime}\in X^{\prime}$, we write $\langle x,x^{\prime}\rangle$ for the dual pairing $x^{\prime}(x)$.
+We say that an operator $T\in\mathcal{L}(X,Y)$ is nuclear provided that it admits a representation of the form
 $T=\sum_{j=1}^{\infty}x_{j}^{\prime}\otimes y_{j},$ (2.4)
-for sequences $(x_{j}^{\prime})\subset X^{\prime}$ and $(y_{j})\subset Y$ satisfying $\sum_{j=1}^{\infty}\|x_{j}^{\prime}\|_{X^{\prime}}\|y_{j}\|_{Y}<\infty$.
-Let $x^{\prime}\in X^{\prime},y\in Y$ , we use the notation $x^{\prime}\otimes y$ for the rank-one operator from $X$ to $Y$ defined by
-$x^{\prime}\otimes y(x)=\langle x,x^{\prime}\rangle y,\quad x\in X.$
+for sequences $(x_{j}^{\prime})\subset X^{\prime}$ and $(y_{j})\subset Y$ satisfying $\sum_{j=1}^{\infty}\|x_{j}^{\prime}\|_{X^{\prime}}\|y_{j}\|_{Y}<\infty.$ For $x^{\prime}\in X^{\prime}$ and $y\in Y$, we denote by $x^{\prime}\otimes y$ the rank-one operator from $X$ to $Y$ defined by x’⊗y(x)=⟨x,x’⟩y, x∈X.
 The nuclear norm of $T$ is given by
-$\|T\|_{\mathcal{N}}:=\inf\sum_{j=1}^{\infty}\|x_{j}^{\prime}\|_{X^{\prime}}\|y_{j}\|_{Y},$
-The infimum in this definition ranges over all representations of the type (2.1).
+∥T∥_N:=inf∑_i=1^∞∥x_i^′∥_X’∥y_i∥_Y, where the infimum in this definition is taken over all representations of the type (2.4).
 We write $\mathcal{N}(X,Y)$ for the class of all nuclear operators mapping $X$ into $Y$.
-Endowed with the norm $\|\cdot\|_{\mathcal{N}}$, it becomes a Banach space, and the following estimate holds for any $A\in\mathcal{L}(X_{0},X)$, $T\in\mathcal{N}(X,Y)$, and $B\in\mathcal{L}(Y,Y_{0})$:
-$\|BTA\|_{\mathcal{N}}\leq\|A\|\,\|T\|_{\mathcal{N}}\,\|B\|.$
+Equipped with the norm $\|\cdot\|_{N}$, it becomes a Banach space.
 In the case where $E$ and $F$ coincide, we write $\mathcal{N}(E)$ for $\mathcal{N}(E,E)$.
-We say that a Banach space $E$ has the approximation property provided that for any compact set $K\subset E$ and any $\epsilon>0$, one can find a finite-rank operator $L:E\to E$ satisfying
-$\|x-Lx\|<\epsilon\quad\text{for all }x\in K.$
+A Banach space $E$ is said to have the approximation property if for every compact set $K\subset E$ and every $\varepsilon>0$, there exists a finite-rank operator $L:E\to E$ such that ∥x-Lx∥<ϵfor all x∈K.
 For a Banach space $E$ with the approximation property, the nuclear trace of an operator $T\in\mathcal{N}(E)$ can be defined by
-$\operatorname{Tr}_{E}(T)=\sum_{j=1}^{\infty}\langle x_{j},x_{j}^{\prime}\rangle,$
-where $T=\sum_{j=1}^{\infty}x_{j}^{\prime}\otimes x_{j}$ is any representation as in (2.1).
-It follows from [10] Theorem V.1.2 that the value of the trace does not depend on the particular representation chosen.
-Standard references for the general theory of operator ideals include [5, 6, 10, 17].
+Tr_E(T)=∑_j=1^∞⟨x_j,x_j^′⟩, where $T=\sum_{j=1}^{\infty}x_{j}^{\prime}\otimes x_{j}$ is any representation as in (2.4).
+It follows from [MR1744872] Theorem V.1.2, that the value of the trace does not depend on the particular representation chosen.
+Standard references for the general theory of operator ideals include [MR1209438, MR1342297, MR1744872, MR2300779].
 Definition 2.1.
 Let $X$ be a Banach space, and let $(\Omega,\Sigma,\mu)$ be any finite measure space, if for every $\mu$-continuous vector measure $G:\Sigma\to X$ of bounded variation, we can find a function $g\in L^{1}(\mu,X)$ for which
-$G(E)=\int_{E}g\,d\mu\qquad\text{holds for all }E\in\Sigma,$
-we say $X$ has the Radon-Nikodým property.
-For each nonnegative integer $n$, define
-$e_{n}(z)=\sqrt{\frac{\alpha^{n}}{n!}}z^{n}.$
-It is well known that $\{e_{n}\}_{n=0}^{\infty}$ forms an orthonormal basis for $F_{\alpha}^{2}$.
-For a general exponent $1<p<\infty$, the same system $\{e_{n}\}_{n=0}^{\infty}$ is also a Schauder basis for $F_{\alpha}^{p}$, as can be seen by expanding each $f\in F_{\alpha}^{p}$ in a Taylor series about the origin.
+$G(E)=\int_{\Omega}g\,d\mu\quad\text{holds for all }E\in\Sigma,$
+we say that $X$ has the Radon-Nikodym property.
 Proposition 2.2.
-For $1<p<\infty$, the Fock space $F_{\alpha}^{p}$ is a Banach space that enjoys both the approximation property and the Radon–Nikodým property.
-The approximation property follows from the fact that $F_{\alpha}^{p}$ has a Schauder basis (see [9, Proposition 2.1]).
-The Radon–Nikodým property holds because $F_{\alpha}^{p}$ is reflexive (see [7, Corollary 3.13]).
-∎
+For $1<p<\infty$, the Fock space $F_{\alpha}^{p}$ is a Banach space that enjoys both the approximation property and the Radon-Nikodym property.
+Proof 1.
+The approximation property follows from [MR4107813, Proposition 2.1]. The Radon-Nikodym property holds because $F_{\alpha}^{p}$ is reflexive(see [MR453964, Corollary 3.13]).
 Let $X$ be a Banach space, if $x\in X$, define $x^{\prime\prime}:X^{\prime}\rightarrow\mathbb{C}$ by
@@
 Let X and Y be Banach spaces such that Y is reflexive. The adjoint map $T\rightarrow T^{*}$ is an isometry from $\mathcal{N}(X,Y)$ onto $\mathcal{N}(Y^{\prime},X^{\prime})$.
+Proof 2.
 Let $T=\sum_{j=1}^{\infty}x_{j}^{\prime}\otimes y_{j}\in\mathcal{N}(X,Y)$, where $(x_{j}^{\prime})\in X^{\prime},(y_{j})\in Y$. For any $y^{\prime}\in Y^{{}^{\prime}},x\in X$, we have
-$\displaystyle T^{*}(y^{\prime})(x)$ $\displaystyle=y^{\prime}(Tx)$
-$\displaystyle=y^{\prime}(\sum_{j=1}^{\infty}x_{j}^{\prime}(x)y_{j})$
-$\displaystyle=\sum_{j=1}^{\infty}x_{j}^{\prime}(x)y^{\prime}(y_{j})$
-$\displaystyle=\sum_{j=1}^{\infty}y_{j}^{\prime\prime}(y^{\prime})x_{j}^{\prime}(x).$
-Then $T^{*}=\sum_{j=1}^{\infty}y_{j}^{\prime\prime}\otimes x_{j}^{\prime}$, and $\|T^{*}\|_{\mathcal{N}}\leq\sum_{j=1}^{\infty}\|y_{j}^{\prime\prime}\|_{Y^{\prime\prime}}\|x_{j}^{\prime}\|_{X^{\prime}}=\sum_{j=1}^{\infty}\|{y_{j}\|_{Y}\|x_{j}^{\prime}\|_{X^{\prime}}}$, the inequality $\|T^{*}\|_{\mathcal{N}}\leq\|T\|_{\mathcal{N}}$ follows by taking the infimum over all representations of the form (2.1).
-On the other hand, let $K\in\mathcal{N}(Y^{\prime},X^{\prime})$, since $Y$ is reflexive, there are sequences $(a_{j}^{\prime})\in X^{\prime}$, $(b_{j})\in Y$ such that $K=\sum_{j=1}^{\infty}b_{j}^{\prime\prime}\otimes a_{j}^{\prime}$, let $L=\sum_{j=1}^{\infty}a_{j}^{\prime}\otimes b_{j}\in\mathcal{N}(X,Y)$, It is easy to check that $L^{*}=K$ and $\|L\|_{\mathcal{N}}\leq\|K\|_{\mathcal{N}}$.
-This prove the desire result.
-∎
-For the dual space of compact operator and nuclear operator, we have the following characterization.
+$T^{*}(y^{\prime})(x)$ $=y^{\prime}(Tx)$
+$=y^{\prime}(\sum_{j=1}^{\infty}x_{j}^{\prime}(x)y_{j})$
+$=\sum_{j=1}^{\infty}x_{j}^{\prime}(x)y^{\prime}(y_{j})$
+$=\sum_{j=1}^{\infty}y_{j}^{\prime\prime}(y^{\prime})x_{j}^{\prime}(x).$
+Then $T^{*}=\sum_{j=1}^{\infty}y_{j}^{\prime\prime}\otimes x_{j}^{\prime}$, and $\|T^{*}\|_{\mathcal{N}}\leq\sum_{j=1}^{\infty}\|y_{j}^{\prime\prime}\|_{Y^{\prime\prime}}\|x_{j}^{\prime}\|_{X^{\prime}}=\sum_{j=1}^{\infty}\|y_{j}\|_{Y}\|x_{j}^{\prime}\|_{X^{\prime}}$.
+Taking the infimum over all possible representations (2.4), we obtain $\|T^{*}\|_{\mathcal{N}}\leq\|T\|_{\mathcal{N}}$.
+On the other hand, let $K\in\mathcal{N}(Y^{\prime},X^{\prime})$.
+Since $Y$ is reflexive, there exist sequences $(a_{j}^{\prime})\in X^{\prime}$ and $(b_{j})\in Y$ such that $K=\sum_{j=1}^{\infty}b_{j}^{\prime\prime}\otimes a_{j}^{\prime}$.
+Let $L=\sum_{j=1}^{\infty}a_{j}^{\prime}\otimes b_{j}\in\mathcal{N}(X,Y)$.
+It is easy to check that $L^{*}=K$ and $\|L\|_{\mathcal{N}}\leq\|K\|_{\mathcal{N}}$.
+This proves the desired result.
+For the dual space of compact operators and nuclear operators, we have the following characterization.
 Proposition 2.4.
-([5, Proposition 16.7]) Let $E$ and $F$ be Banach spaces such that either $E^{\prime\prime}$ or $F^{\prime}$ has the approximation property, and either $E^{\prime\prime}$ or $F^{\prime}$ has the Radon–Nikodým property.
-Then the following isometric identities hold:
-$\displaystyle(\mathcal{K}(E,F))^{\prime}$ $\displaystyle=\mathcal{N}\left(E^{\prime},F^{\prime}\right),$
-$\displaystyle(\mathcal{K}(E,F))^{\prime\prime}$ $\displaystyle=\left(\mathcal{N}\left(E^{\prime},F^{\prime}\right)\right)^{\prime}=\mathcal{L}\left(E^{\prime\prime},F^{\prime\prime}\right),$
-where the duality pairings are given by
-$\displaystyle\langle S,T\rangle_{\mathcal{K},\mathcal{N}}$ $\displaystyle=\sum_{n=1}^{\infty}\left\langle S^{*}y_{n}^{\prime},x_{n}^{\prime\prime}\right\rangle_{E^{\prime},E^{\prime\prime}},$
-$\displaystyle\langle T,S\rangle_{\mathcal{N},\mathcal{L}}$ $\displaystyle=\sum_{n=1}^{\infty}\left\langle y_{n}^{\prime},Sx_{n}^{\prime\prime}\right\rangle_{E^{\prime},E^{\prime\prime}},$
-for $T=\sum_{n=1}^{\infty}x_{n}^{\prime\prime}\otimes y_{n}^{\prime}\in\mathcal{N}\left(E^{\prime},F^{\prime}\right)$.
+([MR1209438, Proposition 16.7]) Let $E$ and $F$ be Banach spaces such that $E^{\prime\prime}$ or $F^{\prime}$ has the approximation property and: $E^{\prime\prime}$ or $F^{\prime}$ has the Radon-Nikodým property.
+Then
+$(\mathcal{K}(E,F))^{\prime}$ $=\mathcal{N}\left(E^{\prime},F^{\prime}\right)$ $\text{isometrically},$
+$(\mathcal{K}(E,F))^{\prime\prime}$ $=\left(\mathcal{N}\left(E^{\prime},F^{\prime}\right)\right)^{\prime}=\mathcal{L}\left(E^{\prime\prime},F^{\prime\prime}\right)$ $\text{isometrically}.$
+The duality brackets are given by
+$\langle S,T\rangle_{\mathcal{K},\mathcal{N}}$ $=\sum_{n=1}^{\infty}\left\langle S^{*}y_{n}^{\prime},x_{n}^{\prime\prime}\right\rangle_{E^{\prime},E^{\prime\prime}}$
+$\langle T,S\rangle_{\mathcal{N},\mathcal{L}}$ $=\sum_{n=1}^{\infty}\left\langle y_{n}^{\prime},Sx_{n}^{\prime\prime}\right\rangle_{E^{\prime},E^{\prime\prime}}$
+if $T=\sum_{n=1}^{\infty}x_{n}^{\prime\prime}\otimes y_{n}^{\prime}\in\mathcal{N}\left(E^{\prime},F^{\prime}\right)$.
 Proposition 2.5.
@@
 where the duality is induced by the trace map:
-$\displaystyle\langle S,T\rangle$ $\displaystyle=Tr(ST)=Tr(TS)$
-$\displaystyle\langle A,B\rangle$ $\displaystyle=Tr(AB)=Tr(BA)$
+$\langle S,T\rangle$ $=\mathrm{Tr}(ST)=\mathrm{Tr}(TS)$
+$\langle A,B\rangle$ $=\mathrm{Tr}(AB)=\mathrm{Tr}(BA)$
 where $S\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$, $T\in\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p})$, $A\in\mathcal{K}(F_{\alpha}^{p},F_{\alpha}^{q})$, $B\in\mathcal{N}(F_{\alpha}^{q},F_{\alpha}^{p})$.
+Proof 3.
 We use $E\cong F$ to denote that the two Banach spaces $E$ and $F$ are isometrically isomorphic. Then, we have
-$\displaystyle\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})^{\prime}$ $\displaystyle\cong\mathcal{N}((F_{\alpha}^{q})^{\prime},(F_{\alpha}^{p})^{\prime})^{\prime}$
-$\displaystyle\cong\mathcal{L}((F_{\alpha}^{q})^{\prime\prime},(F_{\alpha}^{p})^{\prime\prime})$
-$\displaystyle\cong\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p}).$
-The first step is obtained from Theorem 2.3, the second step from Proposition 2.4, the last step follows from the reflexivity.
-Let $S=\sum_{j=1}^{\infty}y_{j}\otimes x_{j}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ and $T\in\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p})$, applying Proposition 2.4, we obtain the duality bracket is given by
-$\displaystyle\langle S,T\rangle$ $\displaystyle=\langle S^{*},\tau T\tau^{-1}\rangle_{\mathcal{N},\mathcal{L}}$
-$\displaystyle=\sum_{j=1}^{\infty}\langle y_{j},\tau Tx_{j}\rangle$
-$\displaystyle=\sum_{j=1}^{\infty}\langle Tx_{j},y_{j}\rangle$
-$\displaystyle=Tr(TS)$
-$\displaystyle=Tr(ST).$
-The other proof is similar. ∎
+$\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})^{\prime}$ $\cong\mathcal{N}((F_{\alpha}^{q})^{\prime},(F_{\alpha}^{p})^{\prime})^{\prime}$
+$\cong\mathcal{L}((F_{\alpha}^{q})^{\prime\prime},(F_{\alpha}^{p})^{\prime\prime})$
+$\cong\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p}).$
+The first step is obtained from Theorem 2.3; the second step from Proposition 2.4; and the last step follows from reflexivity.
+Let $S=\sum_{j=1}^{\infty}y_{j}\otimes x_{j}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ and $T\in\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p})$; applying Proposition 2.4, we obtain that the duality bracket is given by
+$\langle S,T\rangle$ $=\langle S^{*},\tau T\tau^{-1}\rangle_{\mathcal{N},\mathcal{L}}$
+$=\sum_{j=1}^{\infty}\langle y_{j},\tau Tx_{j}\rangle$
+$=\sum_{j=1}^{\infty}\langle Tx_{j},y_{j}\rangle$
+$=\mathrm{Tr}(TS)$
+$=\mathrm{Tr}(ST).$
+The other proof is similar.
 For convenience, in the following sections, we write $A\lesssim B$ to mean that $A\leq CB$ for some positive constant $C$, whose exact value is insignificant.
 The relation $A\simeq B$ signifies that $A\lesssim B\lesssim A$.
-3. Nuclear class Toeplitz operator
-In this section, give a complex Borel measure $\mu$, we are going to characterize when the Toeplitz operator $T_{\mu}$ belongs to nuclear class.
-Let $1\leq p\leq\infty$ and $p^{\prime}$ to be the conjugate exponent of $p$.
+3 Nuclear Toeplitz operator
+In this section, given a complex Borel measure $\mu$, we are going to characterize when the Toeplitz operator $T_{\mu}$ belongs to the nuclear class.
+Let $1\leq p\leq\infty$ and $p^{\prime}$ be the conjugate exponent of $p$.
 For $f\in F_{\alpha}^{p^{\prime}}$, we regard $f$ as an element of $(F_{\alpha}^{p})^{\prime}$ under the pairing $\langle\cdot,\cdot\rangle_{F_{\alpha}^{2}}$, that is
@@
   2. (2)
-    If $T\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$, and $p<q$, then $\widetilde{T}(z)\in L^{r}(dA)$ for all $(\frac{1}{p^{\prime}}+\frac{1}{q})^{-1}\leq r\leq\infty$.
+    If $T\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ and $p\leq q$, then $\widetilde{T}(z)\in L^{r}(dA)$ for all $(\frac{1}{p^{\prime}}+\frac{1}{q})^{-1}\leq r\leq\infty$.
   3. (3)
-    If $T\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$, and $q\leq p$, then $\widetilde{T}(z)\in L^{r}(dA)$ for all $1\leq r\leq\infty$.
+    If $T\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ and $p>q$, then $\widetilde{T}(z)\in L^{r}(dA)$ for all $1\leq r\leq\infty$.
 Moreover,
 $\|\widetilde{T}\|_{L^{r}}\lesssim\|T\|_{\mathcal{N}}.$
+Proof 4.
 To prove (1), let $T=\sum_{j=1}^{\infty}f_{j}\otimes g_{j}$, where $\left(g_{j}\right)\subset F_{\alpha}^{s},\left(f_{j}\right)\subset F_{\alpha}^{1}$ be a nuclear representation. Hence
-$\displaystyle\widetilde{T}(z)$ $\displaystyle=\langle Tk_{z},k_{z}\rangle_{F_{\alpha}^{2}}$
-$\displaystyle=\sum_{j=1}^{\infty}\langle k_{z},f_{j}\rangle_{F_{\alpha}^{2}}\langle g_{j},k_{z}\rangle_{F_{\alpha}^{2}}$
-$\displaystyle=\sum_{j=1}^{\infty}\overline{f_{j}(z)}g_{j}(z)e^{-\alpha|z|^{2}}.$
+$\widetilde{T}(z)$ $=\langle Tk_{z},k_{z}\rangle_{F_{\alpha}^{2}}$
+$=\sum_{j=1}^{\infty}\langle k_{z},f_{j}\rangle_{F_{\alpha}^{2}}\langle g_{j},k_{z}\rangle_{F_{\alpha}^{2}}$
+$=\sum_{j=1}^{\infty}\overline{f_{j}(z)}g_{j}(z)e^{-\alpha|z|^{2}}.$
 Let $1\leq r\leq\infty$, we can find $m\geq 1,n\geq s$ such that $\frac{1}{m}+\frac{1}{n}=\frac{1}{r}$.
-It follows from [19, Theorem 2.10], we have $f_{j}\in F_{\alpha}^{1}\subset F_{\alpha}^{m}$, $g_{j}\in F_{\alpha}^{s}\subset F_{\alpha}^{n}$, and
+It follows from [MR2934601, Theorem 2.10], we have $f_{j}\in F_{\alpha}^{1}\subset F_{\alpha}^{m}$, $g_{j}\in F_{\alpha}^{s}\subset F_{\alpha}^{n}$, and
 $\|f_{j}\|_{\alpha,m}\lesssim\|f_{j}\|_{\alpha,1},\|g_{j}\|_{\alpha,n}\lesssim\|g_{j}\|_{\alpha,s}.$
-Then using Hölder’s inequality, we have
-$\displaystyle\|\widetilde{T}\|_{L^{r}}$ $\displaystyle\leq\sum_{j=1}^{\infty}\|{f_{j}(z)}e^{-\frac{\alpha}{2}|z|^{2}}\|_{L^{m}}\|g_{j}(z)e^{-\frac{\alpha}{2}|z|^{2}}\|_{L^{n}}$
-$\displaystyle\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,m}\|g_{j}\|_{\alpha,n}$
-$\displaystyle\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,1}\|g_{j}\|_{\alpha,s}$
-$\displaystyle\lesssim\sum_{j=1}^{\infty}\|{f_{j}}\|_{(f_{\alpha}^{\infty})^{\prime}}\|g_{j}\|_{\alpha,s}.$
-Now, considering the infimum over all possible representations (2.1) yields
+Then by Hölder’s inequality, we have
+$\|\widetilde{T}\|_{L^{r}}$ $\leq\sum_{j=1}^{\infty}\|{f_{j}(z)}e^{-\frac{\alpha}{2}|z|^{2}}\|_{L^{m}}\|g_{j}(z)e^{-\frac{\alpha}{2}|z|^{2}}\|_{L^{n}}$
+$\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,m}\|g_{j}\|_{\alpha,n}$
+$\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,1}\|g_{j}\|_{\alpha,s}$
+$\lesssim\sum_{j=1}^{\infty}\|{f_{j}}\|_{(f_{\alpha}^{\infty})^{\prime}}\|g_{j}\|_{\alpha,s}.$
+Now, taking the infimum over all possible representations (2.4) yields
 $\|\widetilde{T}\|_{L^{r}}\lesssim\|T\|_{\mathcal{N}}<\infty.$
-The proof of (2) and (3) is similar. ∎
+The proof of (2) and (3) is similar.
 Lemma 3.3.
 Let $1\leq p\leq\infty$, $z\rightarrow k_{z}$ is norm-continuous on $F_{\alpha}^{p}$.
+Proof 5.
 Fix $z_{0}\in\mathbb{C}$, suppose $|z-z_{0}|<1$ and $1\leq p<\infty$. Then
@@
 When $p=\infty$, the conclusion follows from the fact $\|f\|_{\alpha,\infty}\leq\|f\|_{\alpha,p}$ for any $1\leq p<\infty$.
-∎
-Lemma 3.4.
-Let X be a Banach space, and let $\sum_{n=1}^{\infty}x_{n}$ be a absolutely convergent series in X with sum A, Then any rearrangement $\sum_{n=1}^{\infty}x_{\pi(n)}$ induced by a permutation $\pi:\mathbb{N}\rightarrow{\mathbb{N}}$ also converges to A.
-Fix $\epsilon>0$. Choose $N\in\mathbb{N}$ such that:
-$\sum_{k=N+1}^{\infty}\|x_{k}\|<\frac{\epsilon}{2}.$
-Let $S_{n}=\sum_{k=1}^{n}x_{k}$. There exists $N_{1}\geq N$ such that for all $m\geq N_{1}$:
-$\left\|\sum_{n=1}^{m}x_{n}-A\right\|<\frac{\epsilon}{2}.$
-Choose $M\in\mathbb{N}$ such that $\{1,\ldots,N_{1}\}\subseteq\{\pi(1),\ldots,\pi(M)\}$. For any $m>M$, Split the rearranged partial sum:
-$S^{\prime}_{m}=\sum_{\begin{subarray}{c}k\leq m\\ \pi(k)\leq N_{1}\end{subarray}}x_{\pi(k)}+\sum_{\begin{subarray}{c}k\leq m\\ \pi(k)>N_{1}\end{subarray}}x_{\pi(k)}.$
-By the triangle inequality:
-$\|S^{\prime}_{m}-A\|\leq\|S_{N_{1}}-A\|+\sum_{k=N_{1}+1}^{\infty}\|x_{k}\|<\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon.$
-since $\epsilon$ is arbitrary. The rearranged series converges to $A$. ∎
-For a fixed positive number $r$, let
-$r\mathbb{Z}^{2}=\{nr+\mathrm{i}mr:n,m\in\mathbb{Z}\}$
-Then, we define
-$S_{r}=\{z=x+\mathrm{i}y:-r/2\leq x<r/2,\;-r/2\leq y<r/2\}.$
-Clearly, we have the disjoint decomposition for $\mathbb{C}$:
-$\mathbb{C}=\bigcup_{w\in r\mathbb{Z}^{2}}(S_{r}+w),$
-Consequently, for every $f\in L^{1}(\mathbb{C},\mathrm{d}\mu)$ we have
-$\int_{\mathbb{C}}f(z)\,\mathrm{d}\mu(z)=\sum_{w\in r\mathbb{Z}^{2}}\;\int_{S_{r}+w}f(z)\,\mathrm{d}\mu(z).$
+For a fixed positive number $r$, let rZ^2={nr+imr:n,m∈Z}. Then, we define S_r={z=x+iy:-r/2≤x<r/2,-r/2≤y<r/2}. Clearly, we have the disjoint decomposition for $\mathbb{C}$: C=⋃_w∈rZ^2(S_r+w),
 Finally, let $\{a_{n}\}$ denote any fixed enumeration of the square lattice $r\mathbb{Z}^{2}$ as a sequence.
 Theorem 3.4.
-Let $\mu$ be a complex Borel measure on $\mathbb{C}$, and $|\mu|(\mathbb{C})<\infty$. Then $T_{\mu}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$. Moreover, we have
+Let $\mu$ be a complex Borel measure on $\mathbb{C}$ with $|\mu|(\mathbb{C})<\infty$.
+Then $T_{\mu}$ belongs to both $\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{s})$ and $\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ for $1\leq p<\infty$ and $1\leq q,s\leq\infty$.
+Moreover,
 $\|T_{\mu}\|_{\mathcal{N}}\leq\frac{\alpha}{\pi}\,|\mu|(\mathbb{C}).$
-Let $z\in\mathbb{C}$, we consider the operator $k_{z}\otimes k_{z}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$, i.e.
+Proof 6.
+We first proof $T_{\mu}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$. Let $z\in\mathbb{C}$, we consider the operator $k_{z}\otimes k_{z}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$, i.e.
 $(k_{z}\otimes k_{z})f=\langle f,k_{z}\rangle_{F_{\alpha}^{2}}k_{z}.$
@@
 […]
-On the one hand, there exists an interger $N_{1}$ such that $r_{N}<\sqrt{2}\delta$ for all $N>N_{1}$. Fix a $N>max(N_{0},N_{1})$, then by Lemma 3.4, for any $N^{\prime}>N$, we have
-$\displaystyle T_{r_{N}}-T_{r_{N^{\prime}}}=\sum_{k=1}^{\infty}\mu(S_{r_{N}}+a_{k}^{(N)})k_{a_{k}^{(N)}}\otimes k_{a_{k}^{(N)}}-\sum_{j=1}^{\infty}\mu(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})k_{a_{j}^{(N^{\prime})}}\otimes k_{a_{j}^{(N^{\prime})}}$
-$\displaystyle=\sum_{k=1}^{\infty}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})](k_{a_{k}^{(N)}}\otimes k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}}\otimes k_{a_{j}^{(N^{\prime})}})$
-$\displaystyle=I_{1}+I_{2}$
+On the one hand, there exists an integer $N_{1}$ such that $r_{N}<\sqrt{2}\delta$ for all $N>N_{1}$.
+Fix an $N>\max(N_{0},N_{1})$.
+For any $N^{\prime}>N$, recall that in a Banach space, any absolutely summable sequence is unconditionally summable [MR1342297, Proposition 1.1].
+We have,
+$T_{r_{N}}-T_{r_{N^{\prime}}}=\sum_{k=1}^{\infty}\mu(S_{r_{N}}+a_{k}^{(N)})k_{a_{k}^{(N)}}\otimes k_{a_{k}^{(N)}}-\sum_{j=1}^{\infty}\mu(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})k_{a_{j}^{(N^{\prime})}}\otimes k_{a_{j}^{(N^{\prime})}}$
+$=\sum_{k=1}^{\infty}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})](k_{a_{k}^{(N)}}\otimes k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}}\otimes k_{a_{j}^{(N^{\prime})}})$
+$=I_{1}+I_{2}$
 where
-$\displaystyle I_{1}=\sum_{k=1}^{\infty}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})](k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\otimes k_{a_{j}^{(N^{\prime})}}$
-$\displaystyle I_{2}=\sum_{k=1}^{\infty}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]k_{a_{k}^{(N)}}\otimes(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}}).$
+$I_{1}=\sum_{k=1}^{\infty}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})](k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\otimes k_{a_{j}^{(N^{\prime})}}$
+$I_{2}=\sum_{k=1}^{\infty}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]k_{a_{k}^{(N)}}\otimes(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}}).$
 Let
@@
 Thus
-$\displaystyle\|\sum_{k\in S_{1}}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]k_{a_{k}^{(N)}}\otimes(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\|_{\mathcal{N}}$
-$\displaystyle\leq\sum_{k\in S_{1}}\sum_{j=1}^{\infty}|\mu|[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]\|(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\|_{\alpha,1}\|k_{a_{j}^{(N^{\prime})}}\|_{(f_{\alpha}^{\infty})^{\prime}}$
-$\displaystyle\leq|\mu|(\mathbb{C})\epsilon.$
+$\|\sum_{k\in S_{1}}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]k_{a_{k}^{(N)}}\otimes(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\|_{\mathcal{N}}$
+$\leq\sum_{k\in S_{1}}\sum_{j=1}^{\infty}|\mu|[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]\|(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\|_{\alpha,1}\|k_{a_{j}^{(N^{\prime})}}\|_{(f_{\alpha}^{\infty})^{\prime}}$
+$\leq|\mu|(\mathbb{C})\epsilon.$
 On the other hand,
-$\displaystyle\|\sum_{k\in S_{2}}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]k_{a_{k}^{(N)}}\otimes(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\|_{\mathcal{N}}$
-$\displaystyle\leq 2|\mu|(\mathbb{C}\setminus\overline{B(0,R)})$
-$\displaystyle\leq 2\epsilon.$
+$\|\sum_{k\in S_{2}}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]k_{a_{k}^{(N)}}\otimes(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\|_{\mathcal{N}}$
+$\leq 2|\mu|(\mathbb{C}\setminus\overline{B(0,R)})$
+$\leq 2\epsilon.$
 Therefore, we have $\|I_{2}\|_{\mathcal{N}}\leq(2+|\mu|{(\mathbb{C}}))\epsilon$. Similarly, we have $\|I_{1}\|_{\mathcal{N}}\leq C\epsilon$.
@@
 Recall that $\widetilde{\mu}(z)=\frac{\alpha}{\pi}\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}d\mu(w)$, we have
-$\displaystyle|\widetilde{T_{r}}(z)-\frac{\pi}{\alpha}\widetilde{\mu}(z)|=|\sum_{k=1}^{\infty}\mu(S_{r}+a_{k})e^{-\alpha|z-a_{k}|^{2}}-\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}d\mu(w)|$
-$\displaystyle=|\sum_{k=1}^{\infty}\int_{S_{r}+a_{k}}e^{-\alpha|z-w|^{2}}-e^{-\alpha|z-a_{k}|^{2}}d\mu(w)|$
-$\displaystyle\leq\sum_{k=1}^{\infty}\int_{S_{r}+a_{k}}|e^{-\alpha|z-w|^{2}}-e^{-\alpha|z-a_{k}|^{2}}|d|\mu|(w).$
+$|\widetilde{T_{r}}(z)-\frac{\pi}{\alpha}\widetilde{\mu}(z)|=|\sum_{k=1}^{\infty}\mu(S_{r}+a_{k})e^{-\alpha|z-a_{k}|^{2}}-\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}d\mu(w)|$
+$=|\sum_{k=1}^{\infty}\int_{S_{r}+a_{k}}e^{-\alpha|z-w|^{2}}-e^{-\alpha|z-a_{k}|^{2}}d\mu(w)|$
+$\leq\sum_{k=1}^{\infty}\int_{S_{r}+a_{k}}|e^{-\alpha|z-w|^{2}}-e^{-\alpha|z-a_{k}|^{2}}|d|\mu|(w).$
 Since $e^{-\alpha|z|^{2}}$ is uniformly continuous on $\mathbb{C}$, for any $\epsilon>0$, there exists a $r_{0}>0$, if $r<r_{0}$, we have
@@
 for every $w\in S_{r}+a_{k}$. Hence, if $r<r_{0}$, we have
-$\displaystyle|\widetilde{T_{r}}(z)-\frac{\pi}{\alpha}\widetilde{\mu}(z)|<\epsilon.$
-This implies $lim_{r\rightarrow 0}\widetilde{T_{r}}(z)=\frac{\pi}{\alpha}\widetilde{\mu}(z)$, hence $\widetilde{T}(z)=\frac{\pi}{\alpha}\widetilde{\mu}(z)$.
-Set $F(z,w)=T_{\mu}K_{z}(w)$ and $G(z,w)=TK_{z}(w)$.
-These functions are analytic with respect to $w$, and are anti-holomorphic with respect to $z$.
-Furthermore, we have $F(z,z)=\frac{\alpha}{\pi}G(z,z)=e^{\alpha|z|^{2}}\widetilde{\mu}(z)$.
-A standard polarization theorem in several complex variables (see [14]) yields $F(z,w)=\frac{\alpha}{\pi}G(z,w)$ for all $z,w\in\mathbb{C}$.
+$|\widetilde{T_{r}}(z)-\frac{\pi}{\alpha}\widetilde{\mu}(z)|<\epsilon.$
+This implies $\lim\limits_{r\rightarrow 0}\widetilde{T_{r}}(z)=\frac{\pi}{\alpha}\widetilde{\mu}(z)$, hence $\widetilde{T}(z)=\frac{\pi}{\alpha}\widetilde{\mu}(z)$.
+Set $F(z,w)=T_{\mu}K_{z}(w)$ and $G(z,w)=TK_{z}(w)$. These functions are analytic with respect to $w$, and are anti-holomorphic with respect to $z$. Furthermore, we have
+$F(z,z)=\frac{\alpha}{\pi}G(z,z)=e^{\alpha|z|^{2}}\widetilde{\mu}(z).$
+A standard polarization theorem in several complex variables (see [MR1846625]) yields $F(z,w)=\frac{\alpha}{\pi}G(z,w)$ for all $z,w\in\mathbb{C}$.
 Then we have $T_{\mu}K_{z}=\frac{\alpha}{\pi}TK_{z}$ for all $z\in\mathbb{C}$.
@@
 $\|T_{\mu}\|_{\mathcal{N}}=\frac{\alpha}{\pi}\lim\limits_{n\rightarrow\infty}\|T_{r_{n}}\|_{\mathcal{N}}\leq\frac{\alpha}{\pi}|\mu|(\mathbb{C}).$
-Recall that $F_{\alpha}^{1}\subset F_{\alpha}^{p}$ for any $p\geq 1$, and the inclusion is continuous. Therefore, we obtain the following corollary.
-Corollary 3.6.
-Let $\mu$ be a complex Borel measure on $\mathbb{C}$, and $|\mu|(\mathbb{C})<\infty$. Then $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$, where $1\leq p<\infty,1\leq q\leq\infty$. Moreover
-$\|T_{\mu}\|_{\mathcal{N}}\lesssim|\mu|(\mathbb{C}).$
-By the hypothesis and Theorem 3.5, we have $T_{\mu}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$.
-Then we can find $(f_{j})$, $(g_{j})\subset F_{\alpha}^{1}$ such that $T_{\mu}=\sum_{j=1}^{\infty}f_{j}\otimes g_{j}$, and $\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,1}\|g_{j}\|_{\alpha,1}<\infty$.
-Since $p^{\prime},q>1$, the embedding from $F_{\alpha}^{1}$ to $F_{\alpha}^{p^{\prime}}$ and $F_{\alpha}^{q}$ is continuous.
-Hence
-$\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,p^{\prime}}\|g_{j}\|_{\alpha,q}\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,1}\|g_{j}\|_{\alpha,1}<\infty.$
-Then we have $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$, and $\|T_{\mu}\|_{\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})}\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,1}\|g_{j}\|_{\alpha,1}$.
-Take the infimum over all possible representations of $T_{\mu}$ on the right-hand side of the inequality, we obtain $\|T_{\mu}\|_{\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})}\lesssim\|T_{\mu}\|_{\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})}\lesssim|\mu|(\mathbb{C})$.
-∎
-In 2020, Fulsche proved that $T_{f}\in\mathcal{N}(F_{\alpha}^{p})$ if $f\in L^{1}(\mathbb{C})$ (see [9, Lemma 2.11]), thus the above corollary can be viewed as a generalization of this result.
-Moreover, if $\mu$ is a positive measure, the below theorem shows that the condition $\mu(\mathbb{C})<\infty$ is also sufficient for $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ when $q\leq p$.
-Theorem 3.7.
-Let $1\leq q\leq p<\infty,1\leq s\leq\infty$, and let $\mu\geq 0$ satisfying (1.2), then the following statements are equivalent
-  1. (1)
-    $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$;
-  2. (2)
-    $T_{\mu}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{s})$;
-  3. (3)
-    $\mu(\mathbb{C})<\infty$.
-Moreover,
-$\|T_{\mu}\|_{\mathcal{N}}\simeq\mu(\mathbb{C}).$ (3.1)
-By a direct calculation, we see that $\widetilde{T_{\mu}}(z)=\widetilde{\mu}(z)$. Since $\mu\geq 0$, we can use Fubini’s Theorem to obtain
-$\displaystyle\|\widetilde{\mu}\|_{L^{1}}$ $\displaystyle=\frac{\alpha}{\pi}\int_{\mathbb{C}}\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}d\mu(w)dA(z)$
-$\displaystyle=\frac{\alpha}{\pi}\int_{\mathbb{C}}\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}dA(z)d\mu(w)$
-$\displaystyle=\mu(\mathbb{C}).$
+Recall that the normalized reproducing kernel $k_{z}$ is a unit vector in $F_{\alpha}^{p}$ for all $1\leq p\leq\infty$, hence the operator $T_{r}$ defined above belongs to both $\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{s})$ and $\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ for the indices under consideration.
+Repeating this argument completes the proof for all cases.
+From the proof of Theorem 3.4, let
+$S=\{T_{\mu}:|\mu|(\mathbb{C})<\infty\},$
+viewed as a subset of $\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$. Then we have the inclusion
+$S\subset\overline{\operatorname{span}}^{\|\cdot\|_{\mathcal{N}}}\{k_{z}\otimes k_{z}:z\in\mathbb{C}\}.$
+In 2020, Fulsche proved that $T_{f}\in\mathcal{N}(F_{\alpha}^{p})$ if $f\in L^{1}(\mathbb{C})$ (see [MR4107813, lemma 2.11], thus the above Theorem can be viewed as a generalization of this result.
+Moreover, if $\mu$ is a positive measure, Theorem 1.1 shows that the condition $\mu(\mathbb{C})<\infty$ is also sufficient for $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ when $q\leq p$.
+Next, we give the proof of Theorem 1.1:
+Proof 7 (Proof of Theorem 1.1).
+A direct calculation shows that $\widetilde{T_{\mu}}(z)=\widetilde{\mu}(z)$. Since $\mu\geq 0$, we can use Fubini’s Theorem to obtain
+$\|\widetilde{\mu}\|_{L^{1}}$ $=\frac{\alpha}{\pi}\int_{\mathbb{C}}\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}d\mu(w)dA(z)$
+$=\frac{\alpha}{\pi}\int_{\mathbb{C}}\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}dA(z)d\mu(w)$
+$=\mu(\mathbb{C}).$
 $(1)\Rightarrow(3)$: If $1\leq q\leq p<\infty$, by Theorem 3.2 (3), we have $\widetilde{T_{\mu}}(z)=\widetilde{\mu}(z)\in L^{1}(\mathbb{C},dA)$, which implies $\mu(\mathbb{C})<\infty$.
-$(3)\Rightarrow(1)$ follows from Corollary 3.6.
+$(3)\Rightarrow(1)$ follows from Theorem 3.4.
 The proof of $(2)\Leftrightarrow(3)$ is similar.
-Estimate (3.1) follows from the equation $\|\widetilde{\mu}\|_{L^{1}}=\mu(\mathbb{C})$, Theorem 3.2 and Theorem 3.5. ∎
-Now, we present an interesting result. This conclusion is not immediately obvious, but it follows directly from our previous results.
-Corollary 3.8.
-Let $f\in L^{1}(\mathbb{C})$, then $\widetilde{f}\in\cap_{r\geq 1}L^{r}(\mathbb{C})$. Moreover, for any $r\geq 1$, we have $\|\widetilde{f}\|_{L^{r}}\lesssim\|f\|_{L^{1}}$.
-It follows from Theorem 3.5, we have $T_{f}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$, then by Theorem 3.2, we have $\widetilde{f}=\widetilde{T_{f}}\in L^{r}(\mathbb{C})$ for all $\frac{1}{2}\leq r\leq\infty$.
-Moreover, applying Theorem 3.2 and Theorem 3.5, we obtain
-$\|\widetilde{f}\|_{L^{r}}\lesssim\|T_{f}\|_{\mathcal{N}}\lesssim\|f\|_{L^{1}}.$
-Let $\overline{F}_{\alpha}^{p}=\{\overline{f}:f\in F_{\alpha}^{p}\}$, suppose $\mu$ is a complex Borel measure and satisfies condition (1.1), we can define small Hankel operator on $K$ by
-$h_{\mu}(f)(z)=\frac{\alpha}{\pi}\int_{\mathbb{C}}K_{\alpha}(w,z)f(w)\mathrm{e}^{-\alpha|w|^{2}}\,d\mu(w),\quad z\in\mathbb{C}.$
-Theorem 3.9.
-Let $\mu$ be a complex Borel measure on $\mathbb{C}$, and $|\mu|(\mathbb{C})<\infty$.
-Then $h_{\mu}\in\mathcal{N}(F_{\alpha}^{p},\overline{F}_{\alpha}^{q})(\mathcal{N}(f_{\alpha}^{\infty},\overline{F}_{\alpha}^{q}))$, where $1\leq p<\infty,1\leq q\leq\infty$.
-Moreover, we have
-$\|h_{\mu}\|_{\mathcal{N}}\lesssim|\mu|(\mathbb{C}).$
-Following the same method as in the proof of Theorem 3.5, we define
-$h_{r}=\sum_{k=1}^{\infty}\mu(S_{r}+a_{k})\,k_{a_{k}}\otimes\overline{k_{a_{k}}}.$
-Taking $r_{n}=\frac{1}{2^{n}}$, we have $h_{r_{n}}\to h_{\mu}$ in the nuclear norm $\|\cdot\|_{\mathcal{N}}$, which establishes the desired result.
-Definition 3.10.
-Let $X$ and $Y$ be Banach spaces. A linear operator $T\colon X\to Y$ is called absolutely summing provided there exists a constant $C>0$ such that for any finite family $\{x_{j}\}_{j=1}^{n}\subset X$,
-$\sum_{j=1}^{n}\|Tx_{j}\|\;\leq\;C\sup_{\|x^{\prime}\|\leq 1}\sum_{j=1}^{n}|x^{\prime}(x_{j})|\;=\;C\sup_{|\varepsilon_{j}|=1}\Big\|\sum_{j=1}^{n}\varepsilon_{j}x_{j}\Big\|.$
-Every nuclear operator is absolutely summing. In the special cases $X=c_{0}$ or $X=\ell_{\infty}$, the two classes coincide [8, Lemma 2.1].
-Corollary 3.11.
-Let $1\leq s\leq\infty$, and let $\mu\geq 0$ satisfying (1.2), then $T_{\mu}:f_{\alpha}^{\infty}\rightarrow F_{\alpha}^{s}$ is absolutely summing if and only if $\mu(\mathbb{C})<\infty$.
-It is known that $f_{\alpha}^{\infty}$ is isomorphic to $c_{0}$(see [15]). Then the conclusion follows from Theorem 3.7. ∎
+Estimate (1.1) follows from the equation $\|\widetilde{\mu}\|_{L^{1}}=\mu(\mathbb{C})$, Theorem 3.2 and Theorem 3.4.
 Remark 3.5.
 Suppose $1\leq p<q,$ and $\mu\geq 0$.
-If $\mu(\mathbb{C})<\infty$, by Corollary 3.6, we have $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$.
+If $\mu(\mathbb{C})<\infty$, by Theorem 3.4, we have $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$.
 However, we cannot use Theorem 3.2 to prove the equivalence between $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ and $\mu(\mathbb{C})<\infty$, since in this case $(\frac{1}{p^{\prime}}+\frac{1}{q})^{-1}>1$.
@@
 Theorem 3.6.
-Let $1\leq p<\infty$, and let $g$ is a entire function. Then
+Let $1\leq p<\infty$, and let $g$ be an entire function. Then
 $\int_{\mathbb{C}}|f(z)\overline{g(z)}|e^{-\alpha|z|^{2}}dA(z)<\infty$
 for all $f\in F_{\alpha}^{p}$ if and only if $g\in F_{\alpha}^{p^{\prime}}$.
+Proof 8.
 $\Rightarrow:$ For any $n\in\mathbb{N}^{+}$, define the truncated function
@@
 $\varphi_{h}(f)=\int_{\mathbb{C}}f(z)\overline{h(z)}e^{-\alpha|z|^{2}}dA(z),\quad f\in F_{\alpha}^{p}.$
-Then we have $\varphi_{g_{n}}\in(F_{\alpha}^{p})^{\prime}$. Fix $f\in F_{\alpha}^{p}$, since $|g_{n}(z)|\leq|g(z)|$, we can use dominated convergence theorem to obtain
+Then we have $\varphi_{g_{n}}\in(F_{\alpha}^{p})^{\prime}$. Fix $f\in F_{\alpha}^{p}$, since $|g_{n}(z)|\leq|g(z)|$, we can use the dominated convergence theorem to obtain
 $\lim_{n\rightarrow\infty}\varphi_{g_{n}}(f)=\varphi_{g}(f).$
-Hence $\sup_{n}|\varphi_{g_{n}}(f)|<\infty$. The Principle of Uniform Boundedness yields
+Hence $\sup_{n}|\varphi_{g_{n}}(f)|<\infty$. It follows from the Principle of Uniform Boundedness that we have
 $\sup_{n}\|\varphi_{g_{n}}\|=C<\infty.$
@@
 It implies that $\varphi_{g}\in(F_{\alpha}^{p})^{\prime}$.
-Since we can identify the dual space $(F_{\alpha}^{p})^{\prime}$ with $F_{\alpha}^{p^{\prime}}$ through the pairing $\langle\cdot,\cdot\rangle_{F_{\alpha}^{2}}$, we can find $g^{\prime}\in F_{\alpha}^{p^{\prime}}$ such that
+Since $(F_{\alpha}^{p})^{\prime}$ can be identified with $F_{\alpha}^{p^{\prime}}$ under the integral pairing $\langle\cdot,\cdot\rangle_{F_{\alpha}^{2}}$, we can find $g^{\prime}\in F_{\alpha}^{p^{\prime}}$ such that
 $\varphi_{g-g^{\prime}}=0.$
@@
 $\int_{\mathbb{C}}|z^{n}||\overline{(g-g^{\prime})(z)}|e^{-\alpha|z|^{2}}dA(z)<\infty.$
-The dominated convergence theorem then gives
-$\displaystyle\varphi_{g-g^{\prime}}(z^{n})$ $\displaystyle=\int_{\mathbb{C}}z^{n}\overline{(g-g^{\prime})(z)}e^{-\alpha|z|^{2}}dA(z)$
-$\displaystyle=lim_{R\rightarrow{\infty}}\int_{|z|<R}z^{n}\overline{(g-g^{\prime})(z)}e^{-\alpha|z|^{2}}dA(z)$
-$\displaystyle=lim_{R\rightarrow{\infty}}\overline{a_{n}}\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z).$
-$\displaystyle=\overline{a_{n}}\int_{\mathbb{C}}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$
-$\displaystyle=0.$
+Then we can use the dominated convergence theorem to obtain
+$\varphi_{g-g^{\prime}}(z^{n})$ $=\int_{\mathbb{C}}z^{n}\overline{(g-g^{\prime})(z)}e^{-\alpha|z|^{2}}dA(z)$
+$=lim_{R\rightarrow{\infty}}\int_{|z|<R}z^{n}\overline{(g-g^{\prime})(z)}e^{-\alpha|z|^{2}}dA(z)$
+$=lim_{R\rightarrow{\infty}}\overline{a_{n}}\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z).$
+$=\overline{a_{n}}\int_{\mathbb{C}}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$
+$=0.$
 It implies that $a_{n}=0$ for all $n\in\mathbb{N}$. Then we have $g=g^{\prime}\in F_{\alpha}^{p^{\prime}}$.
-The other direction follows from Hölder’s inequality. ∎
+The other direction follows from Hölder’s inequality.
 Example 3.7.
-It is well-known that for $1\leq p<q\leq\infty$, we have the (proper) inclusion $F_{\alpha}^{p}\subset F_{\alpha}^{q}$.
+It is well-known that for $1\leq p<q\leq\infty$, we have the (proper) inclusion $F_{\alpha}^{p}\subset F_{\alpha}^{q}$, and the inclusion is proper.
 Theorem 3.6 tells us: We can find $f\in F_{\alpha}^{p^{\prime}}$ and $g\in F_{\alpha}^{q}$ such that
@@
 $a_{n}=\begin{cases}b^{k}\sqrt{\frac{\alpha^{n_{k}}}{(n_{k})!}}\,n_{k}^{\frac{1}{4}-\frac{1}{2p^{\prime}}}\,\epsilon_{n_{k}},&\text{if }n=n_{k}\text{ for some }k,\\ 0,&\text{otherwise}.\end{cases}$
-The membership of $f$ in $F_{\alpha}^{p^{\prime}}$ is verified by computing two sums. For $1<p^{\prime}\leqslant 2$,
+The membership of $f$ in $F_{\alpha}^{p^{\prime}}$ is verified by computing two sums. For $1<p^{\prime}\leq 2$,
 $\sum_{n=0}^{\infty}|a_{n}|^{p^{\prime}}\left(\frac{n!}{\alpha^{n}}\right)^{\frac{p^{\prime}}{2}}n^{-\frac{p^{\prime}}{4}+\frac{1}{2}}=\sum_{k=1}^{\infty}b^{p^{\prime}k}\epsilon_{n_{k}}^{p^{\prime}}\leq\sum_{k=1}^{\infty}\left(\frac{b^{p^{\prime}}}{2^{p^{\prime}}}\right)^{k}<\infty.$
 If $2\leqslant p^{\prime}<\infty$, then
-$\sum_{n=1}^{\infty}\left|a_{n}\right|^{p}\left(\frac{n!}{\alpha^{n}}\right)^{\frac{p}{2}}n^{\frac{p}{4}-\frac{1}{2}}=\sum_{k=1}^{\infty}b^{pk}n_{k}^{\frac{p}{4}-\frac{1}{2}+\frac{p}{4}-\frac{p}{2p^{\prime}}}\epsilon_{n_{k}}^{p}\leq\sum_{k=1}^{\infty}\left(\frac{b^{p}}{2^{p}}\right)^{k}<\infty.$
-The convergence of these series, together with Theorems 1 and 4 in [18], implies $f\in F_{\alpha}^{p^{\prime}}.$
-Note that the coefficients satisfy
-$a_{n}=\begin{cases}b^{k}\sqrt{\frac{\alpha^{n_{k}}}{(n_{k})!}}\,n_{k}^{\frac{1}{4}-\frac{1}{2q^{\prime}}}\,,&\text{if }n=n_{k},\\ 0,&\text{otherwise}.\end{cases}$
-Consequently,
-$a_{n}\neq o\left(\sqrt{\frac{\alpha^{n}}{n!}}n^{\frac{1}{4}-\frac{1}{2q^{\prime}}}\right)\quad(n\rightarrow\infty),$
-which, by Corollary 5 of [18], means $f\in F_{\alpha}^{p^{\prime}}\setminus F_{\alpha}^{q^{\prime}}$.
+$\sum_{n=1}^{\infty}|a_{n}|^{p}\left(\frac{n!}{\alpha^{n}}\right)^{\frac{p}{2}}n^{\frac{p}{4}-\frac{1}{2}}=\sum_{k=1}^{\infty}b^{pk}n_{k}^{\frac{p}{4}-\frac{1}{2}+\frac{p}{4}-\frac{p}{2p^{\prime}}}\epsilon_{n_{k}}^{p}\leq\sum_{k=1}^{\infty}\left(\frac{b^{p}}{2^{p}}\right)^{k}<\infty.$
+The convergence of these series, together with Theorems 1 and 4 in [MR2215157], implies that $f\in F_{\alpha}^{p^{\prime}}$.
+Since $a_{n}\neq o\left(\sqrt{\frac{\alpha^{n}}{n!}}n^{\frac{1}{4}-\frac{1}{2q^{\prime}}}\right)\quad(n\rightarrow\infty),$ it follows from Corollary 5 of [MR2215157] that $f\in F_{\alpha}^{p^{\prime}}\setminus F_{\alpha}^{q^{\prime}}$.
 A function $g$ is constructed analogously. Let
-$b_{m}=\begin{cases}b^{k}\sqrt{\frac{\alpha^{n_{k}}}{(n_{k})!}}\,n_{k}^{\frac{1}{4}-\frac{1}{2q}}\,\epsilon_{n_{k}},&\text{if }m=n_{k},\\ 0,&\text{otherwise},\end{cases}$
+$b_{m}=\begin{cases}b^{k}\sqrt{\dfrac{\alpha^{n_{k}}}{(n_{k})!}}\,n_{k}^{\frac{1}{4}-\frac{1}{2q}}\,\epsilon_{n_{k}},&\text{if }m=n_{k},\\ 0,&\text{otherwise,}\end{cases}$
 and define $g(z)=\sum b_{m}z^{m}.$ It then follows that $g\in F_{\alpha}^{q}\setminus F_{\alpha}^{p}.$
-If $\int_{\mathbb{C}}|f(z){g(z)}|e^{-\alpha|z|^{2}}dA(z)<\infty$, Applying the dominated convergence theorem, we obtain
-$\displaystyle\int_{\mathbb{C}}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}dA(z)$ $\displaystyle=lim_{R\rightarrow{\infty}}\int_{|z|<R}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}dA(z)$
-$\displaystyle=\lim_{R\rightarrow{\infty}}\sum_{n=0}^{\infty}\sum_{m=0}^{\infty}a_{n}\overline{b_{m}}\int_{|z|<R}z^{n}\overline{z}^{m}e^{-\alpha|z|^{2}}dA(z)$
-$\displaystyle=\lim_{R\rightarrow{\infty}}\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z).$
-From the construction above, we have $a_{n}\overline{b_{n}}\geq 0$ and $\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$ is monotonically increasing in $R$. The monotone convergence theorem yields:
-$\displaystyle\int_{\mathbb{C}}f(z)\overline{g(z)}e^{\alpha|z|^{2}}dA(z)$ $\displaystyle=lim_{R\rightarrow{\infty}}\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$
-$\displaystyle=\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}\int_{\mathbb{C}}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$
-$\displaystyle=\frac{\pi}{\alpha}\sum_{k=1}^{\infty}b^{2k}\epsilon_{n_{k}}$
-$\displaystyle\geq\frac{\pi}{\alpha}\sum_{k=1}^{\infty}(\frac{b^{2}}{3})^{k}$
-$\displaystyle=\infty.$
+If $\int_{\mathbb{C}}|f(z)g(z)|e^{-\alpha|z|^{2}}\mathrm{d}A(z)<\infty,$ then by the dominated convergence theorem, we obtain
+$\int_{\mathbb{C}}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}\mathrm{d}A(z)$ $=\lim_{R\rightarrow\infty}\int_{|z|<R}f(z)\overline{g(z)}e^{-\alpha|z|^{2}}\mathrm{d}A(z)$
+$=\lim_{R\rightarrow\infty}\sum_{n=0}^{\infty}\sum_{m=0}^{\infty}a_{n}\overline{b_{m}}\int_{|z|<R}z^{n}\overline{z}^{m}e^{-\alpha|z|^{2}}\mathrm{d}A(z)$
+$=\lim_{R\rightarrow\infty}\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}\mathrm{d}A(z).$
+From the construction above, we have $a_{n}\overline{b_{n}}\geq 0$ and $\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}\mathrm{d}A(z)$ is monotonically increasing in $R$.
+The monotone convergence theorem yields:
+$\int_{\mathbb{C}}f(z)\overline{g(z)}e^{\alpha|z|^{2}}dA(z)$ $=lim_{R\rightarrow{\infty}}\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}\int_{|z|<R}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$
+$=\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}\int_{\mathbb{C}}|z|^{2n}e^{-\alpha|z|^{2}}dA(z)$
+$=\frac{\pi}{\alpha}\sum_{k=1}^{\infty}\gamma^{2k}\delta_{n_{k}}$
+$\geq\frac{\pi}{\alpha}\sum_{k=1}^{\infty}(\frac{\gamma^{2}}{3})^{k}$
+$=\infty.$
 But $\int_{\mathbb{C}}f(z)\overline{g(z)}\mathrm{e}^{-\alpha|z|^{2}}dA(z)\leq\int_{\mathbb{C}}|f(z){g(z)}|\mathrm{e}^{-\alpha|z|^{2}}dA(z)$. We obtain a contradiction.
-However, for $1\leq p<q\leq\infty$ and $\mu\geq 0$, it remains unknown whether $\mu(\mathbb{C})<\infty$ is necessary for $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$.
 Finally, we conclude this section with an open problem.
-Question 3.15.
+Question 3.8.
 For $1\leq p<q\leq\infty$ and $\mu\geq 0$, what is the necessary and sufficient condition for $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$?
-4. Trace formula and density theorem
+4 Trace formula and density theorem
 In this section, for an operator $T\in\mathcal{N}(F_{\alpha}^{p})$, $(1<p<\infty)$, we will provide some formulas for calculating its trace. Finally, we will establish a density theorem.
@@
 Let $1<p<\infty$ and $T\in\mathcal{N}(F^{p}_{\alpha})$, then we have:
-$Tr(T)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{T}(z)dA(z).$
-Let $T=\sum_{j=1}^{\infty}f_{j}\otimes g_{j}$, where $\left(g_{j}\right)\subset F_{\alpha}^{p},\left(f_{j}\right)\subset F_{\alpha}^{p^{\prime}}$ be a nuclear representation of $T$.
-Then we have $Tr(T)=\sum_{j=1}^{\infty}\langle g_{j},f_{j}\rangle=\sum_{j=1}^{\infty}\langle g_{j},f_{j}\rangle_{F_{\alpha}^{2}}$.
-And the trace is independent of the representation choice.
+$\mathrm{Tr}(T)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{T}(z)dA(z).$
+Proof 9.
+Let $T=\sum_{j=1}^{\infty}f_{j}\otimes g_{j}$ be a nuclear representation of $T$, where $\left(g_{j}\right)\subset F_{\alpha}^{p},\left(f_{j}\right)\subset F_{\alpha}^{p^{\prime}}$.
+Then we have $\mathrm{Tr}(T)=\sum_{j=1}^{\infty}\langle g_{j},f_{j}\rangle=\sum_{j=1}^{\infty}\langle g_{j},f_{j}\rangle_{F_{\alpha}^{2}}$.
+And the trace is independent of the choice of representation.
 Then we have:
-$\displaystyle\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{T}(z)dA(z)$ $\displaystyle=\frac{\alpha}{\pi}\int_{\mathbb{C}}\langle Tk_{z},k_{z}\rangle_{F_{\alpha}^{2}}dA(z)$
-$\displaystyle=\frac{\alpha}{\pi}\int_{\mathbb{C}}\sum_{j=1}^{\infty}\langle K_{z},f_{j}\rangle_{F_{\alpha}^{2}}\langle g_{j},K_{z}\rangle_{F_{\alpha}^{2}}e^{-\alpha|z|^{2}}dA(z)$
-$\displaystyle=\sum_{j=1}^{\infty}\int_{\mathbb{C}}\overline{f_{j}(z)}g_{j}(z)d\lambda_{\alpha}(z)$
-$\displaystyle=\sum_{j=1}^{\infty}\langle g_{j},f_{j}\rangle_{F_{\alpha}^{2}}$
-$\displaystyle=Tr(T).$
-Proposition 4.2.
-Suppose $\mu\geq 0$ and $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p})$, where $1<p<\infty$, then $Tr(T_{\mu})=\frac{\alpha}{\pi}\mu(\mathbb{C})$.
-Let $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p})$, then we have $\mu(\mathbb{C})<\infty$ . By Fubini’s theorem, we obtain:
-$Tr(T_{\mu})$ $=\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{\mu}(z)dA(z)$
-$=(\frac{\alpha}{\pi})^{2}\int_{\mathbb{C}}dA(z)\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}d\mu(w)$
-$=(\frac{\alpha}{\pi})^{2}\int_{\mathbb{C}}d\mu(w)\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}dA(z)$
-$=\frac{\alpha}{\pi}\mu(\mathbb{C}).$
-We now prove a useful result.
-The technical hypotheses in its statement are needed to justify the application of Fubini’s theorem.
-Notice that the case $p=q=2$ of the following theorem appears in [3, Theorem 6], but we adopt a different approach here.
+$\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{T}(z)dA(z)$ $=\frac{\alpha}{\pi}\int_{\mathbb{C}}\langle Tk_{z},k_{z}\rangle_{F_{\alpha}^{2}}dA(z)$
+$=\frac{\alpha}{\pi}\int_{\mathbb{C}}\sum_{j=1}^{\infty}\langle K_{z},f_{j}\rangle_{F_{\alpha}^{2}}\langle g_{j},K_{z}\rangle_{F_{\alpha}^{2}}e^{-\alpha|z|^{2}}dA(z)$
+$=\sum_{j=1}^{\infty}\int_{\mathbb{C}}\overline{f_{j}(z)}g_{j}(z)d\lambda_{\alpha}(z)$
+$=\mathrm{Tr}(T).$
+Notice that the case $p=q=2$ of the following theorem appears in [MR1277446, Theorem 6], but we adopt a different approach here.
 Theorem 4.2.
@@
 Then we have
-$Tr(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(z)\widetilde{S}(z)dA(z).$
+$\mathrm{Tr}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(z)\widetilde{S}(z)dA(z).$
+Proof 10.
 Let $f\in F_{\alpha}^{p}$, we can find a sequence $f_{n}\in K$ such that $f_{n}$ converges to $f$ with $\|\|_{\alpha,p}$.
-The set $K$ is defined as (1.3).
+The ser $K$ is defined as (2.2).
 By assumption (2), we have $T_{\varphi}f_{n}\stackrel{{\scriptstyle\|\cdot\|_{\alpha,q}}}{{\longrightarrow}}T_{\varphi}f$, hence $T_{\varphi}f_{n}(z)\rightarrow T_{\varphi}f(z)$ for every $z\in\mathbb{C}$.
@@
 for all $n\in\mathbb{N}$ and $z\in\mathbb{C}$. Then fix $z\in\mathbb{C}$, by assumption (1), we have
-$\displaystyle|\int_{\mathbb{C}}(f_{n}(w)-f(w))\varphi(w)K(z,w)d\lambda_{\alpha}(w)|$
-$\displaystyle\leq\frac{\alpha}{\pi}\int_{\mathbb{C}}|(f_{n}(w)-f(w))e^{-\frac{\alpha}{2}|z|^{2}}||\varphi(w)K(z,w)e^{-\frac{\alpha}{2}|z|^{2}}|dA(w)$
-$\displaystyle\lesssim\|f_{n}-f\|_{\alpha,p}\|\varphi K_{z}\|_{\alpha,p^{\prime}}\rightarrow 0(n\rightarrow\infty).$
+$|\int_{\mathbb{C}}(f_{n}(w)-f(w))\varphi(w)K(z,w)d\lambda_{\alpha}(w)|$
+$\leq\frac{\alpha}{\pi}\int_{\mathbb{C}}|(f_{n}(w)-f(w))e^{-\frac{\alpha}{2}|z|^{2}}||\varphi(w)K(z,w)e^{-\frac{\alpha}{2}|z|^{2}}|dA(w)$
+$\leq C\|f_{n}-f\|_{\alpha,p}\|\varphi K_{z}\|_{\alpha,p^{\prime}}\rightarrow 0(n\rightarrow\infty).$
 Hence
@@
 for every $f\in F_{\alpha}^{p}$. Then we can use Fubini’s Theorem to obtain
-$Tr(T_{\varphi}S)$ $=\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{T_{\varphi}S}(z)dA(z)$ (4.1)
+$\mathrm{Tr}(T_{\varphi}S)$ $=\frac{\alpha}{\pi}\int_{\mathbb{C}}\widetilde{T_{\varphi}S}(z)dA(z)$ (4.1)
 $=\int_{\mathbb{C}}\langle T_{\varphi}SK_{z},K_{z}\rangle_{F_{\alpha}^{2}}d\lambda_{\alpha}(z)$
@@
 for all $f\in F_{\alpha}^{q},g\in F_{\alpha}^{p^{\prime}}$ . Hence:
-$\displaystyle\int_{\mathbb{C}}K_{u}(z)SK_{z}(u)d\lambda_{\alpha}(z)$ $\displaystyle=\int_{\mathbb{C}}K_{u}(z)\langle SK_{z},K_{u}\rangle_{F_{\alpha}^{2}}d\lambda_{\alpha}(z)$
-$\displaystyle=\int_{\mathbb{C}}K_{u}(z)\langle K_{z},S^{\prime}K_{u}\rangle_{F_{\alpha}^{2}}d\lambda_{\alpha}(z)$
-$\displaystyle=\overline{\int_{\mathbb{C}}K_{z}(u)S^{\prime}K_{u}(z)d\lambda_{\alpha}(z)}$
-$\displaystyle=\overline{S^{\prime}K_{u}(u)}$
-$\displaystyle=SK_{u}(u)$
-$\displaystyle=e^{\alpha|u|^{2}}\widetilde{S}(u).$
+$\int_{\mathbb{C}}K_{u}(z)SK_{z}(u)d\lambda_{\alpha}(z)$ $=\int_{\mathbb{C}}K_{u}(z)\langle SK_{z},K_{u}\rangle_{F_{\alpha}^{2}}d\lambda_{\alpha}(z)$
+$=\int_{\mathbb{C}}K_{u}(z)\langle K_{z},S^{\prime}K_{u}\rangle_{F_{\alpha}^{2}}d\lambda_{\alpha}(z)$
+$=\overline{\int_{\mathbb{C}}K_{z}(u)S^{\prime}K_{u}(z)d\lambda_{\alpha}(z)}$
+$=\overline{S^{\prime}K_{u}(u)}$
+$=e^{\alpha|u|^{2}}\widetilde{S}(u).$
 It follows from (4.1), we have:
-$Tr_{F_{\alpha}^{q}}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(u)\widetilde{S}(u)dA(u).$
-∎ We can now prove
+$\mathrm{Tr}_{F_{\alpha}^{q}}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(u)\widetilde{S}(u)dA(u).$
+We can now prove
 Lemma 4.3.
-Let $1<p,q<\infty$, suppose $\varphi$ is a bounded, compactly supported function on $\mathbb{C}$.
+Let $1<p,q<\infty$, and let $\varphi$ be a bounded, compactly supported function on $\mathbb{C}$.
 Then $T_{\varphi}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$.
-Furthermore, for any $S\in\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p})$ we have:
-$Tr_{F_{\alpha}^{q}}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(z)\widetilde{S}(z)dA(z).$ (4.2)
-$T_{\varphi}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ follows from Corollary 3.6. Applying the pointwise estimates on the Fock space, we have
+Furthermore, for any $S\in\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p})$, we have:
+$\mathrm{Tr}_{F_{\alpha}^{q}}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(z)\widetilde{S}(z)dA(z).$ (4.2)
+Proof 11.
+$T_{\varphi}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ follows from Theorem 3.4. Applying the pointwise estimates on the Fock space, we have
 $|SK_{z}(u)|\leq\|SK_{z}\|_{\alpha,p}e^{\frac{\alpha}{2}|u|^{2}}\leq\|S\|e^{\frac{\alpha}{2}|z|^{2}}e^{\frac{\alpha}{2}|u|^{2}}.$
 Hence
-$\displaystyle\int_{\mathbb{C}}|\varphi(u)|d\lambda_{\alpha}(u)\int_{\mathbb{C}}|K_{u}(z)||SK_{z}(u)|d\lambda_{\alpha}(z)$
-$\displaystyle\lesssim\int_{\mathbb{C}}|\varphi(u)|e^{\frac{\alpha}{2}|u|^{2}}d\lambda_{\alpha}(u)\int_{\mathbb{C}}|K_{u}(z)|e^{-\frac{\alpha}{2}|z|^{2}}dA(z)$
-$\displaystyle\simeq\int_{\mathbb{C}}|\varphi(u)|dA(u)$
-$\displaystyle<\infty.$
-Then the conclusion follows form Theorem 4.3.
-The theorem below extends a result in [3, Theorem 9].
-The cited theorem asserts that the collection of all Toeplitz operators $T_{\varphi}$ with compactly supported continuous symbols is both trace-norm dense in the trace ideal of $F_{\alpha}^{2}$ and operator-norm dense in $\mathcal{K}(F_{\alpha}^{2})$.
-Theorem 4.5.
-Let $\mathcal{C}$ denote the set of all Toeplitz operators $T_{\varphi}$ with continuous, compactly supported symbols $\varphi$ on $\mathbb{C}$, and let $1<p,q<\infty$. Then:
-  1. (1)
-    $\mathcal{C}$ is dense in $\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ with respect to the nuclear norm.
-  2. (2)
-    $\mathcal{C}$ is norm-dense in $\mathcal{K}(F_{\alpha}^{p},F_{\alpha}^{q})$.
+$\int_{\mathbb{C}}|\varphi(u)|d\lambda_{\alpha}(u)\int_{\mathbb{C}}|K_{u}(z)||SK_{z}(u)|d\lambda_{\alpha}(z)$
+$\leq C\int_{\mathbb{C}}|\varphi(u)|e^{\frac{\alpha}{2}|u|^{2}}d\lambda_{\alpha}(u)\int_{\mathbb{C}}|K_{u}(z)|e^{-\frac{\alpha}{2}|z|^{2}}dA(z)$
+$=C\int_{\mathbb{C}}|\varphi(u)|dA(u)$
+$<\infty.$
+Then the conclusion follows from Theorem 4.2.
+We now present the proof of Theorem 1.2:
+Proof 12 (Proof of Theorem 1.2).
 We first prove (2).
-Assume, for contradiction, that $\mathcal{C}$ is not norm-dense in $\mathcal{K}(F_{\alpha}^{p},F_{\alpha}^{q})$.
-An application of the Hahn–Banach theorem combined with Corollary 2.5 implies the existence of a nonzero operator $S\in\mathcal{N}(F_{\alpha}^{q},F_{\alpha}^{p})$ satisfying
-$\operatorname{Tr}(T_{\varphi}S)=0\qquad\text{for all }T_{\varphi}\in\mathcal{C}.$
+Assume, for contradiction, that $C$ is not norm-dense in $\mathcal{K}(F_{\alpha}^{p},F_{\alpha}^{q})$.
+An application of the Hahn-Banach theorem combined with Proposition 2.5 implies the existence of a nonzero operator $S\in\mathcal{N}(F_{\alpha}^{q},F_{\alpha}^{p})$ satisfying
+$\mathrm{Tr}(T_{\varphi}S)=0\qquad\text{for all }T_{\varphi}\in C.$
 By Lemma 4.3, this trace can be written as
-$\operatorname{Tr}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(z)\,\widetilde{S}(z)\,dA(z).$
+$\mathrm{Tr}(T_{\varphi}S)=\frac{\alpha}{\pi}\int_{\mathbb{C}}\varphi(z)\,\widetilde{S}(z)\,dA(z).$
 Since the equality holds for every continuous, compactly supported $\varphi$, it follows that $\widetilde{S}(z)=0$.
@@
 This establishes statement (2).
-The proof of (1) follows a similar argument. ∎
-Funding
-Y. Lu was supported by NNSFC (Grant No.12031002), C. Zu was supported by NNSFC (Grant No.12401151), and the Postdoctoral Researcher Foundation of China (Grant No.GZB20240100).
-Data availability statement
-No data were collected, generated or consulted in connection with this research.
-Declarations
-Conflict of interest
-The authors have no conflict of interest.
+The proof of (1) follows by a similar argument.